### v1

# Learning Task-Agnostic Skill Bases to Uncover Motor Primitives in Animal Behavior

Abstract
Animals flexibly recombine a finite set of core motor primitives to meet diverse task demands, but existing behavior-segmentation methods oversimplify this process by imposing discrete syllables under restrictive generative assumptions. To reflect the animal behavior generation procedure, we introduce skill-based imitation learning (SKIL) for behavior understanding, a reinforcement learning-based imitation framework that (1) infers interpretable skill sets, i.e., latent basis functions of behavior, by leveraging representation learning on transition probabilities, and (2) parameterizes policies as dynamic mixtures of these skills. We validate our approach on a simple grid world, a discrete labyrinth, and unconstrained videos of freely moving animals. Across tasks, it identifies reusable skill components, learns continuously evolving compositional policies, and generates realistic trajectories beyond the capabilities of traditional discrete models. By exploiting generative behavior modeling with compositional representations, our method offers a concise, principled account of how complex animal behaviors emerge from dynamic combinations of fundamental motor primitives.
1 Introduction
Animal behavior is composed of recurring patterns that reflect underlying motor primitives like back grooming, running, and sniffing. Behavior segmentation methods [1, 2, 3, 4, 5] seek to uncover structured patterns in behavior by dividing continuous pose trajectories into discrete units known as stereotyped behavioral syllables. These segmented units can then be used to build structured representations of movement for downstream neurobehavioral analysis.
However, existing behavior segmentation methods face three major limitations. First, they rely on the assumption that behavior consists of discrete action syllables, which oversimplifies the inherently continuous nature of movement and introduces ambiguity during action transitions. Second, they often abstract actions into high-level syllables, overlooking shared motion components across body parts. For example, back and side grooming both involve similar forelimb movements combined with different turning dynamics, and sniffing may occur while walking or sitting, with similar head motion but distinct lower-body patterns. Capturing such compositional structure across body parts is essential for a more detailed understanding of natural behavior. Third, most segmentation models are either non-generative (e.g., clustering [4, 5]) or rely on restrictive generative assumptions (e.g., linear dynamics and hidden Markov model [1, 2, 3]), often leading to unrealistic synthesized behaviors.
To address these limitations, we introduce a new perspective: modeling behavior as emerging from an agent following a reinforcement learning (RL) policy. We study behavioral dynamics and motor primitives by inferring the animal’s policy through an RL–based imitation learning framework. It not only enables more realistic behavior generation through RL but also allows us to discover motor primitives as reusable skill sets that drive behavior toward internal goals. By viewing behavior through this lens, we gain a more flexible, generative, and interpretable understanding of motor primitives that go beyond the constraints of discrete segmentation.
Concretely, we first learn interpretable latent representations from behavior trajectories, which we refer to as skill sets. These represent low-level motion patterns that act as modular building blocks of behavior. Such skills can involve different body parts. By flexibly combining these skill sets, animals can generate a wide range of complex, high-level behaviors. For example, face grooming (forepaw-to-face movements) and body grooming (repetitive strokes along the torso) share common grooming-related skills, while also engaging distinct body parts specific to each movement. We leverage existing advances in representation learning in RL [6, 7, 8] to learn skill sets from transition probability functions. Then, we parametrize policies using these skill sets and apply imitation learning [9] to train the policies from animal behavior demonstrations.
Figure 1: A policy for a movement can be seen as a blend of “vocabularies” from a dictionary containing fundamental motor skills.
Our method is referred to as SKIL, short for Skill-Based Imitation Learning. The fine-grained skill representation enables highly flexible behavioral characterization through the policy. Critically, over long behavioral trajectories, we assume an animal draws from a fixed set of core motor skills, regardless of the specific movements performed [10, 11]. The policy captures how the contribution of each skill evolves continuously over time, reflecting dynamic behavioral changes. Due to the spatial specificity of different motor components, some skills may be engaged briefly, while others persist over longer periods. Moreover, multiple skills can be active simultaneously to construct ongoing movement (Fig. 1). This flexible, compositional view of behavior cannot be achieved using traditional discrete behavioral syllables. SKIL thus offers a more nuanced understanding of how complex actions are generated from dynamic combinations of fundamental motor primitives. Such insight enables researchers to test whether specific fine-grained motor primitives depend on certain local neural circuits, offering a more precise understanding of movement control and its underlying neural mechanisms.
In the remainder of the paper, we introduce preliminaries in Sec. 2, which pave the way for our methods for animal behavior understanding in Sec. 3. We review related work in Sec. 4 and demonstrate the advantage of our methods empirically in Sec. 5. Finally, we conclude with a discussion in Sec. 6.
2 Preliminaries
Markov Decision Processes (MDP). To define skills that characterize mouse behavior, we begin by modeling the observed behavioral trajectories within the framework of MDPs. Formally, an MDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,P,\rho,\gamma,H)$, where $\mathcal{S}$ denotes the state space, capturing both the environment and the animal’s condition—for example, as defined by pose keypoints; $\mathcal{A}$ is the action space denoting feasible movements; $r:\mathcal{S}\times\mathcal{A}\to[0,1]$ is a reward function encoding the immediate utility toward an internal goal; $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the transition operator, with $\Delta(\mathcal{S})$ representing distributions over $\mathcal{S}$; $\rho\in\Delta(\mathcal{S})$ is the initial state distribution; $\gamma\in(0,1)$ is a discount factor; and $H$ is the time horizon. A policy $\pi:\mathcal{S}\times[H]\to\Delta(\mathcal{A})$ is a conditional distribution over actions given a state for each time-step. We assume that the animal generates pose trajectories by following such an MDP where the reward function reflects an intrinsic motivation driving behavior. The behavior is governed by a policy that seeks to maximize this internal reward.
Following standard notations, we define the value function $V(s):=\mathbb{E}\left[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the action-value function $Q(s,a)=\mathbb{E}\left[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$, which are the expected discounted cumulative rewards when executing policy $\pi$. From the above definition, we can establish the following Bellman relationship:
$Q_{h}(s,a)=r(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\bigl{[}V_{h+1}(s^{\prime})\bigr{]},\quad V_{h}(s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}\bigl{[}Q_{h}(s,a)\bigr{]}.$ (1)
Imitation Learning. The goal of imitation learning(IL) is to find a policy $\pi$ that mimics the behavior of the animal demonstrations. In the offline IL setting, we cannot interact with the MDP environment to collect samples using policy $\pi$, but can only access a dataset of transitions sampled from the MDP by the expert, $\mathcal{D}=\left\{(s_{i},a_{i},s^{\prime}_{i})|(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),i=1,2,...,N\right\}$, where $\tau^{e}$ is the data distribution of state-action pairs generated by the expert which is the animal in this study.
[…]
$f^{*}=\mathop{\arg\max}_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\log\frac{f(x_{i},c_{i})}{f(x_{i},c_{i})+\sum_{j=1}^{k}f(y_{i,j},c_{i})},$ (2)
where $\{(x_{i},c_{i})\}_{i=1}^{n}$ are observed samples, $\{y_{i,j}\}_{i\in[n],j\in[k]}$ are artificially generated data sampled from a predefined noise distribution $q(y)$. Here, $k$ represents the ratio of negative to positive samples. This objective can be interpreted as a multi-class extension for distinguishing real data from noise.
3 Skill-based Imitation learning (SKIL) for Behavior Understanding
Given the MDP definition, we frame skill discovery from a control-theoretic perspective. In this view, skills are the fundamental components that enable the construction of diverse policies and reward functions, and thus help explain the motivation behind observed behaviors.
Definition 1 (Skill Set).
Given an arbitrary transition kernel $P(\cdot|s,a)$ in a linear (or low-rank) MDP [15, 16, 17], we can express it via a spectral decomposition:
$P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime}),$ (3)
where $\phi:{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$, $\mu:{\mathcal{S}}\rightarrow\mathbb{R}^{d}$, and $q(s^{\prime})$ is some parametrized probability measure supported on the state space $\mathcal{S}$. We define the function $\phi$ as the skill set. The reward function is then parametrized linearly as $r(s,a)=\phi(s,a)^{\top}w$.
Because animal behavior (e.g., navigation, pose dynamics) follows structured, stereotyped patterns that evolve on low-dimensional manifolds, it can be effectively approximated by low-rank MDPs even in the presence of underlying nonlinearities.
Given the skill definition, substituting Eq. (3) and the linearized reward model into the Bellman equation (Eq. (1)), we get:
$\!Q(s,a)$ $=r(s,a)+\gamma\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}$
$=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]=\phi(s,a)^{\top}u,\!$ (4)
where $u=w+\gamma\int V(s^{\prime})\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$. Thus, the action-value function $Q(s,a)$ can be expressed as a linear combination of skill features $\phi(s,a)$, offering a convenient way to link skills to the policy. Following the maximum entropy reinforcement learning framework [18], we assume the animal’s objective is to maximize the expected reward augmented by the policy’s entropy. Under this assumption, the optimal policy $\pi(a|s)$ can be shown to follow:
$\pi(a|s)=\arg\max_{\pi}\;\;\left[\mathbb{E}_{\pi}[Q(s,a)]+H(\pi)\right]=\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\phi(s,a^{\prime})^{\top}u)},$ (5)
where $H(\pi):=\sum_{a\in\mathcal{A}}\pi(a|s)\,\log\bigl{(}\pi(a|s)\bigr{)}$ is the entropy.
Thereby, the skill function $\phi(s,a)$ not only spans the space of $Q$-functions but also provides a universal basis for constructing the policy $\pi(a|s)$ via Eq. (5). Since $\phi(s,a)$ is derived solely from the transition dynamics $P(s^{\prime}|s,a)$, it remains independent of any specific reward function or task. In this sense, $\phi(s,a)$ encodes intrinsic, general-purpose motor skills available to the animal, while the weight vector $u$ captures task-specific modulations required to produce behavior consistent with different goals. With the skill set established as a foundational representation for both value functions and policies, we are equipped to interpret behavioral trajectories through the lens of skill alignment.
Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (skill discovery), as well as $u$ (skill-based policy learning) from state-action-next-state observations. The learning procedure differs depending on the nature of the behavior data.
3.1 Discrete Version
Skill discovery. For discrete state-action spaces, we apply spectral methods such as singular value decomposition (SVD) [19, 20, 21] or spectral decomposition representation [7, 22] to learn the representations $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$. The resulting skill set $\phi(s,a)$ is then used in the subsequent policy learning stage.
Skill-based policy learning. We now learn the policy $\pi(a|s)$, parameterized by Eq. (5), using maximum likelihood estimation (MLE), i.e., by optimizing the following objective to solve for $u$:
$\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\pi(a|s)\right]=\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\phi(s,a^{\prime})^{\top}u\right)}\right].$ (6)
Limitations for continuous data. While learning from discrete data is relatively straightforward, the continuous case presents additional challenges for two main reasons. First, in the skill discovery step (Eq. (3)), the decomposition $P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ is too restrictive to capture the complexity of continuous behavioral dynamics, such as pose transitions in freely moving animals [2]. For example, consider a common and biologically plausible behavioral model: $s^{\prime}=h(s,a)+\epsilon$, where $h$ is a dynamics function and $\epsilon$ is Gaussian noise. This additive structure, widely used in behavioral modeling, contrasts with the multiplicative form $\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, suggesting that parameterizations preserving additive relationships between $\{s,a\}$ and $s^{\prime}$ are more suitable. Second, in skill-based policy learning, for discrete datasets with a small action space, the denominator (partition function) in Eq. (6) is easy to compute. But for continuous data, the action space is infinite, making it infeasible to enumerate all actions and integrate. In light of these challenges, we adopt an alternative approach to learn the skill representations and policy in continuous state-action spaces.
3.2 Continuous Version
Skill discovery. We model the transition kernel as an energy-based model (EBM) [8]:
[…]
where $\psi(s,a):{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{g}$ and $\nu(s^{\prime}):{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}^{g}$. The goal is to learn $\psi(s,a)$ and $\nu(s^{\prime})$. $Z(s,a)$ is an intractable partition function. This parametrization provides smooth, normalized probabilities with stable gradients, enabling more effective and generalizable representation learning than the unnormalized inner-product model (Eq. (3)).
Proposition 1 (Connection to Skill Definition).
Given the energy-based transition model in Eq. (7), the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
Appendix. A contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
To learn $\psi$ and $\nu$, we sidestep the intractable computation of the partition function $Z(s,a)$ in Eq. (7) by employing NCE (Eq. (2)) to learn the representations.
$\min_{\psi,\,\nu}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),\\ s^{\prime\prime}_{i}\sim\rho\end{subarray}}\left[\log\frac{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))}{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))+\Sigma_{i=1}^{k}\exp(\psi(s,a)^{\top}\nu(s^{\prime\prime}_{i}))}\right],$ (8)
$s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples.
Connection to behavior dynamics. With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right)$. By enforcing unit-norm constraints $\|\psi(s,a)\|^{2}=\|\nu(s^{\prime})\|^{2}=1$, assuming $Z(s,a)$ as a constant and $q(s^{\prime})$ as uniform distribution, as well as taking $\nu$ to be the identity map, we obtain a generalized Gaussian form: $P(s^{\prime}|s,a)=\frac{1}{Z}\exp\left(-\|s^{\prime}-\psi(s,a)\|^{2}\right)$, which aligns with commonly adopted assumptions in animal behavior modeling discussed earlier. Thus, Eq. (7) offers a more general framework that extends traditional dynamics models for studying behavior. Moreover, in practice, using Eq. (7) results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data. Thus, while Eq. (3) is theoretically valid in continuous domains, we adopt the parameterization in Eq. (7) for continuous state and action spaces, as it more effectively supports the learning of skill representations.
Skill-based policy learning. After we obtain the representation $\psi(s,a)$ and $\nu(s^{\prime})$ from Eq. (8), we could theoretically get skill sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix. A). However, since in practice $Z(s,a)$ remains intractable, even with the optimal $\psi$ it is still hard to obtain $\phi$ exactly. Therefore, we introduce a mapping $f:\psi\rightarrow\phi$, parameterized with a lightweight neural network, and learn it via policy learning. Our aim is to learn a function $f$ so that $\phi=f(\psi)$ yields optimal basis functions of policy that best account for the animal behavior data. By applying $\phi=f(\psi)$ to Eq. (3), we obtain $Q(s,a)=f\bigl{(}\psi(s,a)\bigr{)}^{\top}u$.
As mentioned earlier, learning both $f$ and $u$ using the MLE objective in Eq. (6) becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space. This brings us back to the challenge of estimating an unnormalized energy function, $\pi(a|s)\propto\exp(Q(s,a))\propto\exp(f(\psi(s,a))^{\top}u)$. Thus, it’s reasonable to apply NCE here again,
$\min_{f,\,u}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},\\ (s_{i}^{\prime},a^{\prime}_{i})\sim\tau^{e}\end{subarray}}\left[\log\frac{\exp(f(\psi(s,a))^{\top}u)}{\exp(f(\psi(s,a))^{\top}u)+\Sigma_{i=1}^{k}\exp(f(\psi(s,a^{\prime}_{i}))^{\top}u)}\right],$ (9)
where $\{s,a\}$ are positive samples and $\{s^{\prime}_{i},a^{\prime}_{i}\}$ are negative samples.
3.3 Understanding Animal Behavior via Skill and Policy Learning
In this section, we discuss how motor skills $\phi$ and skill weights $u$ can be used to describe animal behavior trajectories, particularly allowing $u(t)$ to vary across tasks or time points. These time-varying coefficients dynamically modulate the influence of each skill. We consider two behavioral modeling scenarios: (1) discrete state-action spaces in a multi-task setting, and (2) continuous state-action spaces in a time-varying setting.
Scenario (1): Consider a mouse navigating a maze [23], where trajectories are discretized into a finite state space (locations) and actions are discrete (up, down, left, right, stay). We assume the animal switches between different reward functions to guide its navigation, with the timing of each reward condition known [24]. We first learn shared skills $\phi(s,a)$ using Eq. (3), then fit task-specific policies $\pi(a|s,t)$ using Eq. (5), where each task $t$ corresponds to one of $T$ reward conditions. This yields $T$ sets of weights $u(t)$, one per task, while sharing a common skill representation across tasks.
Scenario (2): A representative case is a freely behaving mouse [1, 2], where the state is defined by pose keypoints and the action by state change—both continuous. We first learn $\psi(s,a)$ from pose trajectories using Eq. (8). Assuming the policy evolves smoothly over time, we learn $u(t)$ and $f$ via Eq. (9), yielding the motor skill $\phi(s,a)=f(\psi(s,a))$ and time-varying policy $\pi(a|s,t)$. Unlike models with abrupt discrete switches, this continuous-time formulation captures gradual behavioral changes more faithfully over long pose sequences.
Beyond capturing smoothly time-varying skill compositions, our framework allows multiple skills to be active simultaneously. Since the policy is defined as $\pi(a|s,t)\propto\exp\bigl{(}\phi(s,a)^{\top}u(t)\bigr{)}$, each action is a linear combination of basis skills weighted by $u(t)$, enabling overlapping and composable behaviors. For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side—executed concurrently by different body parts. Unlike discrete switching-state models, which assign one behavior per state, our continuous skill-based approach provides a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a fully compositional and continuously time-varying description of animal trajectories.
3.4 Reward Recovery
After estimating $u(t)$ as the skill weights, we can further infer $w(t)$ in the linearized reward model $r(s,a,t)=\phi(s,a)^{\top}w(t)$ as $w(t)=u(t)-\gamma\int V(s^{\prime},t)\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$, where $V(s,t)=\log\sum_{a}\exp Q(s,a,t)$. This allows us to recover the time-varying reward function $r(s,a,t)$ used by animals. Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) [25], where both the policy and underlying reward are inferred. In the context of animal behavior [24, 26, 27], identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal cognition and decision-making processes. Since $V(s,t)$ can be easily computed in closed form in discrete settings, we validate our method by visualizing the inferred rewards in the first two experiments, where the state-action space is finite.
4 Related work
Skill learning from human behavior. Recent work on skill learning for human behavior understanding has primarily focused on robotics, using deep RL and imitation learning to extract reusable primitives from demonstrations [28, 29, 30, 31]. Parallel advances in IRL aim to improve policy inference using low-dimensional bases [32]. However, these approaches are designed for artificial agents in structured settings and do not capture the fine-grained, time-varying structure of naturalistic behavior. In contrast, our work uses RL-based representation learning on real animal data to uncover task-agnostic motor skills and learn a skill-based policy that models continuous, compositional behavior dynamics—capabilities absent in prior work.
Motor primitives learning from animal behavior. Most existing methods for understanding animal behavior identify motor primitives using classification [33, 34], clustering [4, 5], or behavior segmentation techniques [1, 2, 3]. Existing methods label each time point with discrete syllables (e.g., walking, running) and analyze neural correlates of syllable transitions or structure, but treat behavior as rigid and predefined. They fail to capture its compositional, continuously time-varying nature. To our knowledge, no prior work models motor primitives from this flexible perspective. Our approach introduces representation and imitation learning into animal behavior research, enabling dynamic and interpretable modeling of behavior and internal motivation.
5 Experiments
5.1 Application to simulated data in a multi-task gridworld
We evaluate our algorithm on a multi-task gridworld environment. The gridworld consists of a $3\times 3$ lattice with nine discrete states, and each state allows four possible actions: Up, Down, Left, and Right (Fig. 2A). In task $i$, a high reward is assigned to the $(s,a)$ pairs that move toward the location $i$. Fig. 2C (left) shows the ground truth of reward functions for all nine tasks. In each episode, the agent starts from a random start state and must navigate to the task-specific location $i$.
To generate the dataset, we follow this procedure: 1) Use soft value iteration to compute the ground truth Q-function for each task: $Q(s,a,t)=r(s,a,t)+\log\sum_{a}\exp V(s,t)$; 2) Use the resulting Q-function to define the policy: $\pi(a|s,t)=\frac{\exp(Q(s,a,t))}{\sum_{a^{\prime}}\exp(Q(s,a^{\prime},t))}$ and sample trajectories accordingly.
As described in Sec. 3.1, we then learn a set of latent skills and use them to construct the task-specific policy $\pi(a|s,t)$ for each task $t\in\{1,...,9\}$. We assume a total of 64 skills and learn the 64 $\phi$ vectors. Visualizations of the raw features for these skills are shown in the Appendix. B. Using the learned skills, we recover the policy and further infer the reward function used by the agent, as described in Sec. 3.4, with the form: $r(s,a,t)=\phi(s,a)^{\top}w(t)$. This recovered reward (Fig. 2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned skills are sufficient for accurately reconstructing the reward function from behavior data.
To better interpret the learned skills and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig. 2B). Therefore, we visualize the top 8 PC skills and their corresponding task-specific coefficients $w(t)$ in Fig. 2D, treating them as basis vectors spanning the skill space.
The PC skills exhibit interpretable patterns. For instance, in skill 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values. This skill corresponds to moving away from the middle-right grid toward the bottom-left. Examining the $w$ matrix, we see that skill 1 has a strong positive weight for task 8, consistent with the goal of moving toward the bottom-right corner in that task. In contrast, skill 0 contributes negatively to task 8, as it promotes movement toward the bottom-left and away from the goal. Similar interpretations can be made for other skills and tasks.
Overall, in this gridworld example, we successfully recover reward functions from behavior trajectories, and importantly, the learned $\phi$ and $w$ are both interpretable and aligned with task goals.
Figure 2: A. State-action map: each of the nine grids is divided into four cells representing action values (Up, Down, Left, Right). In task $i$, high reward is assigned to $(s,a)$ pairs moving toward the $i$th location. B. Explained variance of the top 15 principal components; variance drops near zero after PC7. C. Left: true rewards for all 9 tasks (yellow = high, green = low). Right: recovered rewards. D. Reward weight $w$ and top 8 PC skills from the $\phi$ matrix. Reward weights indicate the contribution of the top 8 PC skills to each task. In the PC skill plot, red indicates positive feature values, blue indicates negative.
5.2 Application to animal navigation behavior
5.2.1 Dataset and model setup
We next evaluate our algorithm on real animal behavior data from [23]. In this experiment, a thirsty mouse is trained to navigate in a maze (Fig. 3A), starting each trial from the central home cage and attempting to reach a water port located at the end of one branch. Although the mouse’s behavior is primarily driven by water foraging after training, it also exhibits exploration of unvisited areas and returns to the home cage for shelter. This complex behavior cannot be captured by a single reward function. The state space has the structure of a binary tree and is defined by the mouse’s location. The actions include moving to its left parent, right parent, left child, and right child. Non-existence of the moving target will end in staying still.
We first apply the EM-based segmentation algorithm from [24] to divide the behavior into three interpretable tasks: water seeking, home seeking, and exploration. The algorithm could also infer reward maps for these tasks, which are shown in Fig. 3A (top row). While the mouse’s true internal reward functions remain unknown, we treat these inferred rewards as effective ground truth since they generate policies that closely replicate the observed behavior. Thus, we are able to segment long navigation trajectories into task-specific episodes, yielding a multi-task dataset with discrete $(s,a)$ pairs, analogous to our simulated gridworld setting.
To uncover a task-agnostic set of skills capable of constructing all three reward functions, we apply our model as described in Sec. 3.1. The learned $\phi$ vectors serve as a policy basis, encoding movement tendencies such as approaching or avoiding certain regions of the maze.
Figure 3: A. True and recovered rewards for the three tasks. B. Two dominant skills for the water and home tasks respectively. Each skill indicates that taking a specific action (orange arrow) toward the dark purple state yields a high value. C. $w$ values for all tasks, colored by task; blue stars highlight high-weight skills for water seeking, and orange for home seeking.
5.2.2 Results
To assess model performance, we estimate the task-specific weights $w(t)$ and reconstruct the reward function as $r(s,a,t)=\phi(s,a)^{\top}w(t)$. The recovered reward functions align closely with the ground truth (Fig. 3A, bottom row). In the water-seeking task, the recovered reward has peaks near the water port and along the path leading to it. In the home-seeking task, a distinct peak appears at the home cage. In the exploration task, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
We further visualize the learned skills by applying PCA to obtain PC skills. See the Appendix. C for raw features of skills. Fig. 3B displays two top-contributing skills for the water and home tasks respectively, selected based on the peak of linear weights $w$ (Fig. 3C). The water-related skills promote movement toward the water port, while the home-related skills guide navigation to the cage. Notably, the skills important for one task have minimal or negative contributions to the other (Fig. 3C), indicating clear functional specialization.
These results show that our model accurately recovers multiple reward functions from real behavior and, importantly, learns interpretable skills whose contributions to each task are both distinct and behaviorally meaningful.
5.3 Application to animal free-moving behavior
5.3.1 Dataset and model setup
In this section, we apply our method to a dataset of free-moving mouse behaviors to extract motor skills and analyze pose dynamics, specifically how complex pose sequences are dynamically composed of underlying motor primitives. The dataset [2] contains the keypoint coordinates of eight mouse body parts, including four points along the spine, the head, the nose, and both ears. Each dimension of the state corresponds to either $x$ or $y$ coordinate of a body part. The action at each timestep is defined as the velocity of the state, computed as $a_{t}=(s_{t+1}-s_{t})/\delta t$. We set the time interval to $\delta t=1$ ($1/30s$ in the original dataset). We formulate the data using a continuous MDP.
As outlined in Sec. 3.2, we parameterize the transition function as $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})\right)$, and use NCE to learn the feature functions $\psi(s,a)$ and $\nu(s^{\prime})$. We then apply a transformation function $f$ and again use NCE to learn the skill representation $\phi(s,a)=f(\psi(s,a))$ and the time-dependent weights $u(t)$. To ensure temporal smoothness in the learned weights $u(t)$, we place a Gaussian random walk prior over the trajectories: $u(t)\sim\mathcal{N}(u(t-1),\sigma^{2}I)$. We optimize $f$ and $u(t)$ using coordinate descent, updating them alternately. For this dataset, our focus is on understanding the learned policy structure, which reflects the dynamics of animal poses. Therefore, we do not perform IRL in this setting and instead concentrate on interpreting the learned skill representation $\phi$ and the temporal weights $u(t)$.
We split the full dataset into training and test trajectories in an 8:2 ratio. We first learn both $f$ and $u(t)$ on the training set. Then, given the learned $f$, we estimate $u(t)$ on the test set. Here, $f$ is a time-invariant model parameter shared across all time, while $u(t)$ is a time-dependent variable that must be inferred separately for each test trajectory and cannot be transferred from training.
We compare our SKIL method with a baseline model, Keypoint-MoSeq [2], a state-of-the-art behavior segmentation algorithm that partitions continuous pose trajectories into discrete behavioral syllables. For a fair comparison, we train Keypoint-MoSeq on the training set to learn its parameters and then use it to infer behavioral syllables, i.e., perform segmentation, on the test trajectories.
5.3.2 Results
We evaluate performance using the area under the Receiver Operating Characteristic (ROC) curve (AUC), which quantifies the model’s ability to distinguish positive from negative samples. AUC is chosen because it allows direct comparison between our unnormalized energy function and Keypoint-MoSeq’s likelihood score. Given $(s,a),(s^{\prime},a^{\prime})\sim\tau^{e}$, we define positive samples as $(s,a)$ and negative samples as mismatch pairs $(s,a^{\prime})$. We use the log-likelihood of the autoregressive model in Keypoint-MoSeq as its prediction score and the negative energy function $f(\psi(s,a))^{\top}u(t)$ as our prediction score, respectively.
We repeat the experiment 10 times and report the results as a box plot in Fig. 4A. Our model achieves higher AUC scores on both training and test sets (paired t-test, $p_{train}=0.0037,p_{test}=0.0046$), demonstrating a stronger ability to distinguish positive from negative samples. This suggests that SKIL more accurately captures time-varying pose dynamics through smoothly evolving motor primitives, while Keypoint-MoSeq’s reliances on discrete switching syllables produce a coarser representation of the underlying complexity.
Beyond quantitative comparisons, we also qualitatively visualize and interpret the key skills $\phi(s,a)$ associated with example pose dynamics. We examine the time-varying weights $u(t)$ and compare them with the behavioral syllable segmentation produced by Keypoint-MoSeq (Fig. 4B). We present five example clips of observed animal behavior, each consisting of five consecutive frames. Each clip is characterized by a unique combination of motor skills. For each clip, we show the top 1-2 most dominant skills. For each skill, we display the animal’s skeleton with red arrows showing the motion field, computed by averaging the actions that most strongly activate that skill. The movement semantics of each skill are labeled above the visualizations. A more comprehensive visualization of all learned skills is included in the Appendix. D.
By combining the discovered skills with real animal behavior, we assess the interpretability of each skill in the five clips (Fig. 4B). First, the right-turn behavior in clip 1 is captured by the dominance of two rightward skills (skill 1 and skill 2). A stronger movement at the head is reflected by a higher value of skill 1. Clips 2, 3, and 4 show the mouse turning right, pausing to groom its head and ears, and then continuing to turn right. The alternate dominance of skill 3 and skill 4 aligns well with the behavior dynamics. Clip 5 shows a simultaneous behavioral mixture of moving forward and sniffing, and is captured by the equal strength of skill 5 and skill 6. Across all skills, skill 4 appears across clips 2, 3, and 4, showing its general utility. The transitions and mixtures of behaviors are effectively reflected in the learned skills and their temporal weights $u(t)$.
However, when we examine the segmentation produced by Keypoint-MoSeq and the syllable dynamics (Fig. 4C), we find inconsistencies. For example, clear rightward turning in clip 1 is barely visible in syllable 3 to which it is assigned. Clips 2 and 3 are both assigned to syllable 2, even though clip 2 shows pure turning right while clip 3 is dominated by grooming movements. For clip 5, the mixture of fast moving forward and sniffing is not reflected in syllable 2.
Taken together, these results show that compared with similar approaches, SKIL provides a more accurate interpretation of pose dynamics and could capture more complex behaviors through a compact, task-agnostic set of motor skills. This offers us a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motor primitives.
Figure 4: A. AUC on training and test sets. B. We present example learned skills, behavior clips, $u(t)$, and segmentation from Keypoint-MoSeq. C. Syllables identified by Keypoint-MoSeq (adopted from [2]).
6 Discussion
We introduce the SKIL framework that advances animal behavior modeling by offering a flexible, interpretable alternative to traditional discrete behavioral segmentation methods. It learns task-agnostic motor primitives (skill sets) that dynamically compose into complex behaviors. SKIL effectively recovers meaningful skills, reconstructs reward functions, and captures nuanced behavioral dynamics, providing a principled and quantitative model of how complex policies arise from fundamental motor components.
Several limitations remain to be addressed in future work. First, the accuracy of inferred skills is sensitive to input data quality, as occlusions or tracking errors can degrade performance. Additionally, training the energy-based models can also be computationally intensive, potentially restricting real-time or large-scale applications. Lastly, while the framework uncovers abstract motor primitives, establishing direct correspondences between these learned "skills" and specific neural circuit dynamics still requires further experimental validation.
Beyond advancing animal behavior research, SKIL has broader implications. Positively, a better understanding of motor control mechanisms could, for instance, inform new treatments for movement disorders or inspire more adaptable AI. On the other side, extending these principles to model human behavior carries ethical risks, such as perpetuating or amplifying societal biases present in training data. A robust ethical framework is essential to mitigate such risks in the development and application of these technologies.
Appendix A Approximating Energy-Based Formulation with Low-Rank Spectral Decomposition
By simple algebra, we obtain the quadratic potential function,
$P(s^{\prime}|s,a)$ $\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right).$ (10)
The term $\exp\left(-\frac{\|\psi(s,a)-\nu(s^{\prime})\|^{2}}{2}\right)$ is the Gaussian kernel, for which we apply the random Fourier feature [6, 35] and obtain the spectral decomposition of Eq. (7) as
$P(s^{\prime}|s,a)=\langle\phi_{\omega}(s,a),\mu_{\omega}(s^{\prime})q(s^{\prime})\rangle_{\mathcal{N}(\omega)},$ (11)
where $\omega\sim\mathcal{N}(0,I)$ is the frequency in the Fourier domain, and
$\phi_{\omega}(s,a)$ $=\exp\left(-i\omega^{\top}\psi(s,a)\right)\exp\left(\|\psi(s,a)\|^{2}/2-\log Z(s,a)\right),$ (12)
$\mu_{\omega}(s^{\prime})$ $=\exp\left(-i\omega^{\top}\nu(s^{\prime})\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right).$ (13)
Note that Eq. (11) needs infinite $\omega$ to calculate the expectation. To connect it to finite dimension $\phi(s,a)\in\mathbb{R}^{d},\mu(s^{\prime})\in\mathbb{R}^{d}$, we use the Monte-Carlo method to approximate it with finite samples,
$P(s^{\prime}|s,a)\approx\frac{1}{M}\sum_{i=1}^{M}\phi_{\omega_{i}}(s,a)\mu_{\omega_{i}}(s^{\prime})q(s^{\prime}).$ (14)
Introduce vectors $\phi(s,a)$ and $\mu(s^{\prime})$ such that
$\phi(s,a)$ $:=\frac{1}{\sqrt{M}}\left[\phi_{\omega_{1}}(s,a),\phi_{\omega_{2}}(s,a),...,\phi_{\omega_{M}}(s,a)\right],$ (15)
$\mu(s^{\prime})$ $:=\frac{1}{\sqrt{M}}\left[\mu_{\omega_{1}}(s^{\prime}),\mu_{\omega_{2}}(s^{\prime}),...,\mu_{\omega_{M}}(s^{\prime})\right].$ (16)
Then it’s straightforward to see that,
$\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})=\frac{1}{M}\sum_{i=1}^{M}(\phi_{\omega_{i}}(s,a)^{\top}\mu_{\omega_{i}}(s^{\prime}))\approx P(s^{\prime}|s,a).$ (17)
Hence, Eq. (7) can, in principle, yield the skill representation introduced earlier.
Appendix B Learned Skills in the Multi-task Gridworld Dataset
Figure 5: A. State-action maps for all 64 skills. B. Reward weight $w$ for all 64 skills.
We visualize all original 64 skills (Fig. 5) introduced in Sec. 5.1. It shows some meaningful patterns as mentioned before. For example, skill 0 assigns high values to those $(s,a)$ pairs leading to the middle-middle grid and the bottom-middle grid, and assigns low values to the up-left grid and the up-right grid. Thus, it is employed negatively in Task 0 (up-left reward) and Task 2 (up-right reward). However, Task 4 (middle-middle) didn’t use this skill and used skill 39 negatively instead. The complex many-to-many relationship between skills and tasks informs us of the redundancy in the original skills, which inspires us to use PCA to analyze the principal components of the skill space and simplify the skill weights. It could be seen from the comparison between Fig. 2D and Fig. 5A that principal components are a less redundant description of the skill space.
Appendix C Learned Skills in the Animal Navigation Behavior Dataset
Figure 6: A. Left. The value for each $(s,a)$ pair in skill 0 before PCA. Right. The most important $(s,a)$ pairs. B. Left. The value for each $(s,a)$ pair in skill 0 after PCA. Right. The most important $(s,a)$ pairs. C. Boxplot for the effective dimensions before and after PCA.
In the original skills of the labyrinth environment, multiple $(s,a)$ pairs are simultaneously activated, so it is rather hard to analyze which $(s,a)$ pairs are the most important ones that could represent the focus and function of the skill. Given the redundancy of the skill sets, as in Appendix. B, we perform PCA to analyze the principle components of the skill space and simplify the skill representation. To show the effect of PCA, we plot one skill (skill 0) before (Fig. 6A right) and after PCA (Fig. 6B right). Basically, we only want to show the most important pairs in one map and do not want low-value pairs to disturb the visualization. To determine how many $(s,a)$ pairs are important, we sort the $(s,a)$ pairs based on the value $\phi(s,a)$ in skill 0, i.e., the first feature of the output of $\phi(s,a)$ (Fig. 6A left and B left). It could be seen straightforwardly that after PCA, the skill becomes more concentrated on several $(s,a)$ pairs. We calculate the mean $\mu$ and variance $\sigma$ across all skills and all dimensions and take $\mu+\sigma$ as the threshold, above which $(s,a)$ pairs are deemed the most important ones and are shown on the right. We show 80 pairs before PCA and 8 pairs after PCA. The number of the most important pairs in each skill is called the “effective dimension.” The effective dimension is calculated across all skills (Fig. 6C). Paired t-test ($p=1.3\times 10^{-51}$) shows that there exists a significant decrease of effective dimensions after PCA. So the map becomes more distinct and functionally separated.
Appendix D Learned Skills in the Animal Free-moving Behavior Dataset
Figure 7: A. Policy weight $u(t)$ for all 64 skills. B. The motion field for all 64 skills learned from the video, computed by averaging the states and actions that most strongly activate each skill.
We show all skills learned from the 200-timestep video clip of the free-moving mouse mentioned in Sec. 5.3.2. We have completed the visualization of those skills that were previously omitted due to their perceived lack of importance. Due to the increased number of displayed skills, we had to renumber each skill. We show the present skill number above the skill motion field figure, and previous numbers (if applicable) in the parentheses following the present number.