### v2

# Learning Task-Agnostic Motifs to Capture the Continuous Nature of Animal Behavior

Abstract
Animals flexibly recombine a finite set of core motor motifs to meet diverse task demands, but existing behavior segmentation methods oversimplify this process by imposing discrete syllables under restrictive generative assumptions. To better capture the continuous structure of behavior generation, we introduce motif-based continuous dynamics (MCD) discovery, a framework that (1) uncovers interpretable motif sets as latent basis functions of behavior by leveraging representations of behavioral transition structure, and (2) models behavioral dynamics as continuously evolving mixtures of these motifs. We validate MCD on a multi-task gridworld, a labyrinth navigation task, and freely moving animal behavior. Across settings, it identifies reusable motif components, captures continuous compositional dynamics, and generates realistic trajectories beyond the capabilities of traditional discrete segmentation models. By providing a generative account of how complex animal behaviors emerge from dynamic combinations of fundamental motor motifs, our approach advances the quantitative study of natural behavior.
1 Introduction
A critical direction in animal behavior research has been identifying recurring patterns, often referred to as stereotyped behavioral syllables, like back grooming, running, and sniffing, directly from large-scale behavior recordings. Behavior segmentation methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Hsu & Yttri, 2021; Berman et al., 2014) seek to uncover such structured patterns in behavior by dividing continuous pose trajectories into discrete syllables. Classic behavior segmentation approaches can be categorized into three groups: (1) supervised classification (Marks et al., 2022; Segalin et al., 2021), (2) clustering-based methods (Hsu & Yttri, 2021; Berman et al., 2014; Whiteway et al., 2021), and (3) switching-dynamics-based methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Costacurta et al., 2022). The segmented syllables can then be used to build structured representations of movement for downstream neurobehavioral study.
However, existing behavior segmentation methods face four major limitations. First, they rely on the assumption that behavior consists of discrete action syllables, which oversimplifies the inherently continuous nature of movement and introduces ambiguity during action transitions. Second, they often extract actions as abstract syllables that fail to capture how individual body parts contribute to different motions. For example, back and side grooming both involve similar forelimb movements combined with different turning dynamics, and sniffing may occur while walking or sitting, with similar head motion but distinct lower-body patterns. Capturing such compositional structure across body parts is essential for a more detailed understanding of natural behavior. Third, most segmentation models are either non-generative (e.g., clustering (Hsu & Yttri, 2021; Berman et al., 2014)) or rely on restrictive generative assumptions (e.g., linear dynamics and Markov models (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022)), often leading to unrealistic synthesized behaviors. Fourth and most importantly, classification/clustering-based methods ignore temporal dependencies and fail to capture motifs reflecting true pose dynamics, while dynamics-based methods impose linear or nonlinear dynamics assumptions that may not match actual animal behavior, making such assumptions potentially harmful.
To address these limitations, we introduce a new perspective: modeling behavior under the reinforcement learning (RL) framework. We study behavioral dynamics and motor motifs by inferring the animal’s policy through an RL-based imitation learning (IL) framework. It not only enables more realistic behavior generation through RL but also allows us to discover reusable motor motif sets to construct a policy driven by internal rewards. By viewing behavior through this lens, we gain a more flexible, generative, and interpretable understanding of motor motifs that go beyond the constraints of discrete segmentation. Note that Aldarondo et al. (2024) also applied RL-based IL to analyze animal behavior, but without parsing long untrimmed behaviors into fine-grained, interpretable motor motifs, so their work lies outside the scope of behavior segmentation considered here.
We hypothesize that animals draw from a fixed set of core motor motifs to construct diverse movements over long behavioral trajectories (Santuz et al., 2019; Flash & Hochner, 2005). Building on this, we propose Motif-based Continuous Dynamics discovery (MCD) to parse long trajectories and uncover motifs and policies that reflect behavioral dynamics. Concretely, we learn interpretable latent representations, or motif sets, via spectral decomposition–based representation learning in RL (Dai et al., 2014; Ren et al., 2023; Shribak et al., 2024). These motifs correspond to low-level motion patterns serving as modular building blocks of behavior and can involve different body parts. For instance, face grooming (forepaw-to-face) and body grooming (torso strokes) share grooming motifs while engaging distinct body parts. These motifs can then be used to sufficiently represent policies that characterize complex high-level behaviors. Finally, we apply imitation learning to train motif-based policies from demonstrations. This framework leverages RL in two aspects: (1) motifs are inferred through RL-based representation learning, and (2) we use policies formed from motifs to characterize behavioral dynamics. As shown later, both aspects avoid any model assumptions while capturing motifs and policies that faithfully reflect behavioral dynamics.
Figure 1: A policy for a movement can be seen as a blend of “vocabularies” from a dictionary containing fundamental motor motifs.
Another key innovation in constructing policies from motifs is that each motif’s contribution evolves continuously over time, reflecting dynamic behavioral changes. Some motifs may be brief, others prolonged, and multiple motifs can be active simultaneously to build ongoing movement (Fig. 1). This flexible, compositional view cannot be achieved with traditional discrete syllables. MCD thus provides a nuanced account of how complex actions arise from dynamic motif combinations and enables testing whether fine-grained motifs depend on specific neural circuits. Compared with prior segmentation methods, MCD offers soft segmentation that captures continuous time-varying processes rather than discrete switches.
Note that there is a rich literature in robotics on learning skills through imitation learning (Paraschos et al., 2013; Lioutikov et al., 2017; Li et al., 2017; Ajay et al., 2021; Peng et al., 2022; Kuang et al., 2025). We use RL-based imitation not to compete with existing skill learning methods, as most are not suited to the behavior segmentation task for neuroscience study in this paper. Rather, our goal is to employ the RL framework as a principled way to characterize the continuous nature of animal behavior while rendering fine-grained motor motifs, analogous to how Keypoint-MoSeq relies on an SLDS framework (Weinreb et al., 2024) and VAME relies on an autoencoder framework (Luxem et al., 2022). To summarize, we contribute to behavior understanding for neuroscience and neuroethology through the following points:
  • We introduce the first RL-based IL framework for behavior segmentation, a fundamental advance since RL naturally treats behavior as a decision-making process shaped by policies and rewards. Unlike dynamics-based methods, it explains why behaviors occur, not just how they unfold.
  • Within this framework, we propose RL-based representation learning to discover motif-based policies. The learned motifs and policies do not rely on dynamics assumptions or any model assumptions. They can faithfully characterize behavioral dynamics without the mismatch issues of prior segmentation methods.
  • Our method reveals the continuous nature of animal behavior, moving beyond the discrete segmentation assumptions of existing methods.
  • It provides a nuanced understanding of how complex behaviors emerge from dynamic combinations of fundamental motor motifs.
2 Preliminaries
Markov Decision Processes (MDP). To define motifs that characterize animal behaviors, we begin by modeling the observed behavioral trajectories within the framework of MDPs. Formally, an MDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,P,\rho,\gamma,H)$, where $\mathcal{S}$ denotes the state space, capturing both the environment and an animal’s condition—for example, positions of pose keypoints; $\mathcal{A}$ is the action space denoting feasible movements; $r:\mathcal{S}\times\mathcal{A}\to[0,1]$ is a reward function encoding the immediate utility toward an internal goal; $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the transition operator, with $\Delta(\mathcal{S})$ representing distributions over $\mathcal{S}$; $\rho\in\Delta(\mathcal{S})$ is the initial state distribution; $\gamma\in(0,1)$ is a discount factor; and $H$ is the time horizon. A policy $\pi:\mathcal{S}\times[H]\to\Delta(\mathcal{A})$ is a conditional distribution over actions given a state for each time-step. We assume that an animal generates pose trajectories by following such an MDP where the reward function reflects an intrinsic motivation driving behavior. The behavior is governed by a policy that seeks to maximize this internal reward.
Following standard notations, we define the value function $V(s):=\mathbb{E}\left[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the action-value function $Q(s,a)=\mathbb{E}\left[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$, which are the expected discounted cumulative rewards when executing policy $\pi$. From the above definition, we can establish the following Bellman relationship:
$Q_{h}(s,a)=r(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\bigl[V_{h+1}(s^{\prime})\bigr],\quad V_{h}(s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}\bigl[Q_{h}(s,a)\bigr].$ (1)
Offline Imitation Learning. We use imitation learning(IL) to find a policy $\pi$ that mimics animal behavior. In the offline IL setting, we cannot interact with the MDP environment to collect samples using policy $\pi$, but can only access a dataset of transitions sampled from the MDP by the expert, $\mathcal{D}=\left\{(s_{i},a_{i},s^{\prime}_{i})|(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),i=1,2,...,N\right\}$, where $\tau^{e}$ is the data distribution of state-action pairs generated by the expert which is the animal in this study.
3 Motif-based Continuous Dynamics (MCD) Discovery
Given the MDP definition, we frame motif discovery from a control-theoretic perspective. In this view, motifs are the fundamental components that enable the construction of diverse policies and reward functions, and thus help explain the motivation behind observed behaviors.
Definition 1 (Motif Set).
Given an arbitrary transition kernel $P(\cdot|s,a)$ in an MDP, we can express it via a spectral decomposition:
$P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime}),$ (2)
where $\phi:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$, $\mu:\mathcal{S}\rightarrow\mathbb{R}^{d}$, and $q\in\Delta(\mathcal{S})$ is a parametrized probability distribution over the state space. We define the function $\phi$ as the motif set. The reward function is then parametrized linearly as $r(s,a)=\phi(s,a)^{\top}w$.
Spectral decomposition has been widely studied in RL representation learning (Ren et al., 2023; Zhang et al., 2022; Shribak et al., 2024). We adopt this approach here to define motifs given the transition kernel and define rewards accordingly. Since spectral decomposition derives latents directly from the transition kernel without model assumptions, motif learning is thus independent of model assumptions and faithfully reflects the motifs present in the behavior data.
Given the motif definition, substituting Eq. 2 and the linearized reward model into the Bellman equation (Eq. 1), we get:
$Q(s,a)=r(s,a)+\gamma\!\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]=\phi(s,a)^{\top}u,\!$ (3)
where $u=w+\gamma\int V(s^{\prime})\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$. Thus, the action-value function $Q(s,a)$ can be expressed as a linear combination of motif features $\phi(s,a)$, offering a convenient way to link motifs to the policy. Following the maximum entropy reinforcement learning framework (Haarnoja et al., 2018), we assume the animal’s objective is to maximize the expected reward augmented by the policy’s entropy. Under this assumption, the optimal max-entropy policy $\pi(a|s)$ can be shown to follow:
$\pi(a|s)=\arg\max_{\pi}\;\;\left[\mathbb{E}_{\pi}[Q(s,a)]+H(\pi)\right]=\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\phi(s,a^{\prime})^{\top}u)},$ (4)
where $H(\pi):=\sum_{a\in\mathcal{A}}\pi(a|s)\,\log\bigl(\pi(a|s)\bigr)$ is the entropy.
Proposition 1.
The policy in Eq. 4 is not based on any model assumption but emerges naturally as the max-entropy policy based on spectral decomposition of the transition kernel. Furthermore, the learned motifs $\phi$ can represent any max-entropy policy through an appropriate choice of $u$.
The reason we define $\phi(s,a)$ as motifs is that $\phi$ provides the linear basis for the environment transition, policies, and rewards. Policies characterize behavioral dynamics by describing action tendencies conditioned on state (how behaviors evolve). Combined with the environment dynamics $P(s^{\prime}|s,a)$, it induces the transition distribution $P(s^{\prime}|s)=\sum_{a}P(s^{\prime}|s,a)\,\pi(a|s)$, which governs the evolution of behavior trajectories, as is often directly modeled in dynamics-based methods (Wiltschko et al., 2015; Weinreb et al., 2024). Rewards, in turn, reflect the underlying driving factors of these trajectories (why behaviors evolve). Moreover, because $\phi$ is derived solely from the transition dynamics $P(s^{\prime}|s,a)$, it remains independent of any specific reward or task. In this sense, $\phi(s,a)$ encodes intrinsic, general-purpose motor motifs available to animals, while the weight vector $u$ captures task-specific modulations required to produce behavior aligned with different goals. Thus, we can interpret behavioral trajectories through the lens of motifs $\phi$.
From Def. 1 and Prop. 1, we conclude that the learned motifs and policies do not rely on model assumptions, yet the policies faithfully capture behavioral dynamics as action tendencies conditioned on state. Thus, our method is assumption-free while capturing dynamics, unlike classification/clustering methods (no dynamics) or dynamics-based methods (restrictive assumptions).
Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (motif discovery), as well as $u$ (motif-based policy learning that characterizes the continuous behavioral dynamics) from demonstrations. The learning procedure differs depending on the nature of the behavior data (i.e. discrete or continuous).
3.1 Discrete Version
Motif discovery. For discrete state-action spaces, we apply spectral methods such as singular value decomposition (SVD) (Golub & Reinsch, 1971; Golub & Van Loan, 2013; Trefethen & Bau, 2022) or spectral decomposition representation (Ren et al., 2023; HaoChen et al., 2021) to learn the representations $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$. The resulting motif set $\phi(s,a)$ is then used in the subsequent policy learning stage.
Motif-based policy learning. We now learn the policy $\pi(a|s)$, parameterized by Eq. 4, using maximum likelihood estimation (MLE), i.e., by optimizing the following objective to solve for $u$:
$\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\pi(a|s)\right]=\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\phi(s,a^{\prime})^{\top}u\right)}\right].$ (5)
3.2 Continuous Version
While learning from discrete data is relatively straightforward, the continuous case presents additional challenges for two main reasons. First, in the motif discovery step (Eq. 2), the decomposition $P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ is too restrictive to capture the complexity of continuous behavioral dynamics, such as pose transitions in freely moving animals (Weinreb et al., 2024). For example, consider a common and biologically plausible behavioral model: $s^{\prime}=h(s,a)+\epsilon$, where $h$ is a dynamics function and $\epsilon$ is Gaussian noise. This additive structure, widely used in behavioral modeling, contrasts with the multiplicative form $\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, suggesting that parameterizations preserving additive relationships between $\{s,a\}$ and $s^{\prime}$ are more suitable. Second, in motif-based policy learning, for discrete datasets with a small action space, the denominator (partition function) in Eq. 5 is easy to compute. But for continuous data, the action space is infinite, making it infeasible to enumerate all actions and integrate. In light of these challenges, we adopt an alternative approach to learn the motif representations and policy in continuous state-action spaces.
Motif discovery. We model $P(s^{\prime}|s,a)$ as an energy-based model (EBM) (Shribak et al., 2024):
$P(s^{\prime}|s,a)=q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})-\log Z(s,a)\right),Z(s,a)=\int q(s^{\prime})\exp(\psi(s,a)^{\top}\nu(s^{\prime}))ds^{\prime},$ (6)
where $\psi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{g}$ and $\nu:\mathcal{S}\to\mathbb{R}^{g}$ are neural-network feature maps. Here, $Z(s,a)$ is an intractable partition function. Compared to the unnormalized inner-product model (Eq. 2), this EBM formulation yields smooth, normalized probabilities and stable gradients, leading to more effective and generalizable motif representations.
Proposition 2 (Connection to Motif Definition).
Given the EBM model in Eq. 6, the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
Appendix. C contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
To learn $\psi$ and $\nu$, we employ noise-contrastive estimation (NCE) (Ma & Collins, 2018; Gutmann & Hyvärinen, 2010; 2012), which enables optimization of unnormalized statistical models without explicitly computing the partition function. In this way, we sidestep the intractable computation of $Z(s,a)$ in Eq. 6 by solving
$\min_{\psi,\,\nu}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),\\ s^{\prime\prime}_{i}\sim\rho\end{subarray}}\left[\log\frac{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))}{\exp(\psi(s,a)^{\top}\nu(s^{\prime}))+\Sigma_{i=1}^{k}\exp(\psi(s,a)^{\top}\nu(s^{\prime\prime}_{i}))}\right],$ (7)
where $s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples from distribution $\rho$.
Connection to behavioral dynamics. With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right)$ from Eq. 6. By enforcing unit-norm constraints $\|\psi(s,a)\|^{2}=\|\nu(s^{\prime})\|^{2}=1$, assuming $Z(s,a)$ as a constant and $q(s^{\prime})$ as uniform distribution, as well as taking $\nu$ to be the identity map, we obtain a generalized Gaussian form: $P(s^{\prime}|s,a)=\frac{1}{Z}\exp\left(-\|s^{\prime}-\psi(s,a)\|^{2}\right)$, which aligns with commonly adopted assumptions in animal behavior modeling discussed earlier. Thus, Eq. 6 offers a more general framework that extends traditional dynamics models for studying behavior. Moreover, in practice, using Eq. 6 results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data. Thus, while Eq. 2 is theoretically valid in continuous domains, we adopt the parameterization in Eq. 6 for continuous state and action spaces, as it more effectively supports the learning of motif representations underlying animal behavior.
Motif-based policy learning. After we obtain the representation $\psi(s,a)$ and $\nu(s^{\prime})$ from Eq. 7, we could theoretically get motif sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix. C). However, since in practice $Z(s,a)$ remains intractable, even with the optimal $\psi$ it is still hard to obtain $\phi$ exactly. Therefore, we introduce a mapping $f:\psi\rightarrow\phi$, parameterized with a neural network, and learn it via policy learning. Our aim is to learn a function $f$ so that $\phi=f(\psi)$ yields optimal basis functions of policy that best account for the animal behavior data. By applying $\phi=f(\psi)$ to Eq. 3, we obtain $Q(s,a)=f\bigl(\psi(s,a)\bigr)^{\top}u$.
As mentioned earlier, learning both $f$ and $u$ using the MLE objective in Eq. 5 becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space. This brings us back to the challenge of estimating an unnormalized energy function, $\pi(a|s)\propto\exp(Q(s,a))\propto\exp(f(\psi(s,a))^{\top}u)$. Thus, it’s reasonable to apply NCE here again,
$\min_{f,\,u}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},\\ (s_{i}^{\prime},a^{\prime}_{i})\sim\tau^{e}\end{subarray}}\left[\log\frac{\exp(f(\psi(s,a))^{\top}u)}{\exp(f(\psi(s,a))^{\top}u)+\Sigma_{i=1}^{k}\exp(f(\psi(s,a^{\prime}_{i}))^{\top}u)}\right],$ (8)
where $\{s,a\}$ are positive samples and $\{s^{\prime}_{i},a^{\prime}_{i}\}$ are negative samples.
3.3 Understanding Animal Behavior via Motif and Policy Learning
In this section, we discuss how motor motifs $\phi$ and motif weights $u$ can be used to describe animal behavior trajectories, particularly allowing $u$ to vary across tasks $t$ or time points $t$. As the motif coefficients, $u(t)$ would dynamically modulate the influence of each motif on the final policy. We consider two behavioral modeling scenarios: (1) discrete state-action spaces in a multi-task setting, and (2) continuous state-action spaces in a time-varying setting. The concrete results will be later shown in Sec. 4.2 and Sec. 4.3 respectively. In either case, $u(t)$ would be a matrix where each column corresponds to one weight for one task or time point.
Scenario (1): Consider a mouse navigating a maze (Rosenberg et al., 2021), where trajectories are discretized into a finite state space (locations) and actions are discrete (up, down, left, right, stay). We assume the animal switches between $T$ strategies, each associated with a unique reward, that guide its navigation, with the timing of each reward condition known from Ke et al. (2025). We first learn shared motifs $\phi(s,a)$ using Eq. 2, then fit task-specific policies $\pi(a|s,t)$ using Eq. 5, where each task $t$ corresponds to one of $T$ strategies. This yields $T$ sets of weights $u(t)$, one per task, while sharing a common motif set across tasks.
Scenario (2): A representative case is a freely behaving mouse (Wiltschko et al., 2015; Weinreb et al., 2024), where the state is defined by pose keypoints and the action by state change–both continuous. We first learn $\psi(s,a)$ from pose trajectories using Eq. 7. Assuming the policy evolves smoothly over time, we learn $u(t)$ and $f$ via Eq. 8, yielding the motor motif $\phi(s,a)=f(\psi(s,a))$ and time-varying policy $\pi(a|s,t)$. Unlike models with abrupt discrete switches, this continuous-time formulation captures gradual behavioral changes more faithfully over long pose sequences.
Beyond capturing smoothly time-varying motif compositions, our framework allows multiple motifs to be active simultaneously. Since the policy is defined as $\pi(a|s,t)\propto\exp\bigl(\phi(s,a)^{\top}u(t)\bigr)$, each action is a generalized linear combination of basis motifs weighted by $u(t)$, enabling overlapping and composable behaviors. For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side—recruiting different motifs concurrently. Unlike discrete switching-state models, which assign one behavior per state, our continuous motif-based approach provides a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a fully compositional and continuously time-varying description of animal trajectories. In Sec. 4, we would show (1) what motifs we have learned, and (2) how they are used to construct the final policy, on one simulation datasets and two real animal behavior datasets.
3.4 Reward Recovery
After estimating $u(t)$ as the motif weights for policy construction, we can further infer $w(t)$ for reward representation $r(s,a,t)=\phi(s,a)^{\top}w(t)$ as $w(t)=u(t)-\gamma\int V(s^{\prime},t)\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$, where $V(s,t)=\log\sum_{a}\exp Q(s,a,t)$. This allows us to recover the time-varying reward function $r(s,a,t)$ used by animals. Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) (Ziebart et al., 2008), where both the policy and underlying reward are inferred from demonstrations. In the context of animal behavior (Ke et al., 2025; Zhu et al., 2024; Ashwood et al., 2022), identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal cognition and decision-making processes. Since $V(s,t)$ can only be easily computed in closed form in discrete settings, we validate our method by visualizing the inferred rewards in the first two experiments, where the state-action space is discrete and finite.
4 Experiments
4.1 Application to simulated data in a multi-task gridworld
The gridworld consists of a $3\times 3$ lattice with nine discrete states, and each state allows four possible actions: Up, Down, Left, and Right (Fig. 2A). In task $i$, a high reward is assigned to the $(s,a)$ pairs that move toward the location $i$. Fig. 2C (left) shows the ground truth of reward functions for all nine tasks. In each episode, the agent starts from a random start state and must navigate to the task-specific location $i$. See Appendix. D for more details for data generation.
As described in Sec. 3.1, we learn a set of latent motifs and use them to construct the task-specific policy $\pi(a|s,t)$ for each task $t\in\{1,...,9\}$. Because the computational complexity only scales linearly with the number of motifs, we assume a total of 64 motifs and learn the 64 $\phi$ vectors to cover the motif space as much as possible. Visualizations of these motifs are shown in the Appendix. D. Using the learned motifs, we recover the policy and further infer the reward function, as described in Sec. 3.4, with the form: $r(s,a,t)=\phi(s,a)^{\top}w(t)$. This recovered reward (Fig. 2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned motifs are sufficient for accurately reconstructing the reward function from behavior data.
To better interpret the learned motifs and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig. 2B). Therefore, we visualize the top 8 PC motifs and their corresponding task-specific coefficients $w(t)$ in Fig. 2D, treating them as basis vectors spanning the motif space. The PC motifs exhibit interpretable patterns. For instance, in motif 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values. This motif corresponds to moving away from the middle-right grid toward the bottom-left. Examining the $w$ matrix, we see that motif 1 has a strong positive weight for task 8, consistent with the goal of moving toward the bottom-right corner in that task. In contrast, motif 0 contributes negatively to task 8, as it promotes movement toward the bottom-left and away from the goal. Similar interpretations can be made for other motifs and tasks.
In this gridworld experiment, we successfully recover reward functions from behavior trajectories, and, importantly, the learned $\phi$ and $w$ are effectively deployed in different task settings.
Figure 2: A. State-action map: each of the nine grids is divided into four cells representing action values (Up, Down, Left, Right). In task $i$, high reward is assigned to $(s,a)$ pairs moving toward the $i$th location. B. Explained variance of the top 15 principal components; variance drops near zero after PC7. C. Left: true rewards for all 9 tasks (yellow = high, green = low). Right: recovered rewards. D. Reward weight $w$ and top 8 PC motifs from the $\phi$ matrix. Reward weights indicate the contribution of the top 8 PC motifs to each task. In the PC motif plot, red indicates positive feature values, blue indicates negative.
4.2 Application to animal navigation behavior
Dataset and model setup.
We next evaluate our method on a real animal behavior dataset from Rosenberg et al. (2021). In this experiment, a thirsty mouse is trained to navigate in a binary-tree maze (Fig. 3A), starting each trial from the central home cage and attempting to reach a water port at one leaf node. The state space is defined by the mouse’s location on the tree. The actions include moving to its left parent, right parent, left child, and right child. Although the mouse’s behavior is primarily driven by water foraging, it also exhibits exploration of unvisited areas and returns to the home cage for shelter. This complex behavior cannot be captured by a single reward function. Studying this dataset allows us to discover motifs shared across multiple reward functions and policies. This in turn tests whether complex navigation behavior, under multiple competing motivations, can be distilled into a small set of reusable motifs that provide interpretable insight into animal decision-making.
We first apply the segmentation algorithm from Ke et al. (2025) to divide long behavior trajectories into three interpretable tasks: water seeking, home seeking, and exploration. The algorithm could also infer reward maps for these tasks, which are shown in Fig. 3A (top row). While the mouse’s true internal reward functions remain unknown, we treat these inferred rewards as effective ground truth since they generate policies that closely replicate the observed behavior. Thus, we are able to segment long navigation trajectories into task-specific episodes, yielding a multi-task dataset with discrete $(s,a)$ pairs, analogous to our simulated gridworld setting. To uncover a task-agnostic set of motifs capable of constructing all three reward functions, we apply our model as described in Sec. 3.1. Again, to cover the motif space as much as possible, we set the number of motifs $d=128$.
Figure 3: A. True and recovered rewards for the three tasks. B. Two dominant motifs for the water and home tasks respectively. Each motif indicates that taking a specific action (orange arrow) toward the dark purple state yields a high value. C. $w$ values for all tasks, colored by task; blue stars highlight high-weight motifs for water seeking, and orange for home seeking, all shown in (B).
Results.
To assess model performance, we estimate the task-specific weights $w(t)$ and reconstruct the reward function as $r(s,a,t)=\phi(s,a)^{\top}w(t)$. The recovered reward functions align closely with the ground truth (Fig. 3A, bottom row). In the water-seeking task, the recovered reward has peaks near the water port and along the path leading to it. During home-seeking, a distinct peak appears at the home cage. When exploring, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
We further visualize the learned motifs by applying PCA to obtain PC motifs. See the Appendix. E for raw features of motifs. Fig. 3B displays two top-contributing motifs for the water and home tasks respectively, selected based on the peak of linear weights $w$ (Fig. 3C). The water-related motifs promote movement toward the water port, while the home-related motifs guide navigation to the cage. Notably, the motifs important for one task have minimal or negative contributions to the other (Fig. 3C), indicating clear functional specialization. Fig. 3C also shows non-task-specific motifs with similar weights in both water- and home-seeking tasks.
These results show that our model not only recovers multiple reward functions from real behavior but also learns interpretable motifs whose contributions to each task are distinct and behaviorally meaningful. Unlike previous reward discovery on this dataset (Ashwood et al., 2022; Ke et al., 2025), which cannot identify such motifs, our approach reveals how reward maps can be decomposed into smaller, reusable action components. This decomposition offers a mechanistic view of how local decision processes combine to produce strategies such as water-seeking or home-seeking.
4.3 Application to animal free-moving behavior
Dataset and model setup.
We finally apply our method to a continuous dataset of free-moving mouse behaviors (Weinreb et al., 2024) to extract motor motifs and analyze fine-grained pose dynamics. This dataset contains the keypoint coordinates of eight mouse body parts, including the head, the nose, both ears, and four spine nodes. Each dimension of the state corresponds to either $x$ or $y$ coordinate of a body part. The action at each timestep is defined as the velocity of state, $a_{t}=(s_{t+1}-s_{t})/\delta t$. We set the time interval to $\delta t=1$ ($1/30s$ in the original dataset). We formulate the data using a continuous MDP. Studying this dataset allows us to ask whether free-moving behaviors, which often appear as mixtures of grooming, locomotion, and postural adjustments, can be represented as combinations of a compact set of task-agnostic motor motifs.
As outlined in Sec. 3.2, we use NCE to learn the motif representation $\phi(s,a)$ and the time-dependent weights $u(t)$. To ensure temporal smoothness in the learned weights $u(t)$, we place a Gaussian random walk prior over the trajectories: $u(t)\sim\mathcal{N}(u(t-1),\sigma^{2}I)$. To perform more stable optimization, we optimize $\phi(s,a)$ and $u(t)$ using coordinate descent, updating them alternately. For this dataset, our focus is on understanding the learned policy structure, which reflects the dynamics of animal poses. Therefore, we do not perform IRL in this setting and instead concentrate on interpreting the learned motif representation $\phi$ and the temporal weights $u(t)$. With respect to the choice of the number of motifs, we find in practice that the performance grows more slowly once past $d=64$, so we choose this as an optimal number. See Appendix. F for more training details.
We compare our MCD method with two representative behavior segmentation approaches: (1) Keypoint-MoSeq (Weinreb et al., 2024), as a representative for switching-dynamics-based segmentation methods; and (2) SemiSeg (Whiteway et al., 2021), as a representative for clustering-based behavior segmentation methods. Note that Keypoint-MoSeq is regarded as a SOTA approach, because it extends the autoregressive hidden Markov model (AR-HMM) and MoSeq (Wiltschko et al., 2015), and has been shown to outperform B-SOiD (Hsu & Yttri, 2021), VAME (Luxem et al., 2022), and MotionMapper (Berman et al., 2014). We also include OPAL (Ajay et al., 2021), a representative autoencoder-based motif learning algorithm from robotics, as a baseline to highlight the advantages of our method and why robotics approaches are ill-suited for behavioral segmentation in neuroscience.
Results.
We evaluate performance using the area under the Receiver Operating Characteristic (ROC) curve (AUC), which quantifies the model’s ability to distinguish positive from negative samples. AUC is chosen because it allows direct comparison between our unnormalized energy function and Keypoint-MoSeq’s likelihood score. Given $(s,a),(s^{\prime},a^{\prime})\sim\tau^{e}$, we define positive samples as $(s,a)$ and negative samples as mismatch pairs $(s,a^{\prime})$. With respect to the choice of the prediction score, for Keypoint-Moseq, we use the action log-likelihood. For MCD, we use the negative energy function $\phi(s,a)^{\top}u(t)$. For SemiSeg, we assume action variance=1 and use the Gaussian log-likelihood of actions. For OPAL, we use the action log-likelihood.
We repeat the experiment 10 times and report the results as box plots in Fig. 4A. Our model achieves the highest AUC on both training and test sets (paired t-test, $p<0.05$ for every baseline), demonstrating the strongest ability to distinguish positive from negative samples. This suggests that MCD accurately captures time-varying pose dynamics through smoothly evolving motifs, while other models fail. Keypoint-MoSeq’s reliances on discrete switching syllables produce a coarser representation of the underlying complexity, and the autoencoders in the other two models have weaker expressive ability.
Beyond quantitative comparisons, we also qualitatively visualize and interpret the key motifs $\phi(s,a)$ associated with example pose dynamics. For MCD, we examine the time-varying weights $u(t)$ of a long animal behavior video (length=250) and choose five example animal behavior clips (length=5) to check the interpretability of the result (Fig. 4B). For our model, each clip is characterized by a unique combination of motor motifs. For each clip, we show the top 1-2 most dominant motifs. For each motif, we display the animal’s skeleton with red arrows showing the motion field, computed by averaging the actions that most strongly activate that motif. The movement semantics of each motif are labeled above the visualizations. A more comprehensive visualization of all learned motifs is included in the Appendix. F. For comparison, we run other models on the same behavior video, showing the latent representations by SemiSeg (Fig. 4C) and OPAL (Fig. 4D). To show clearer results for SemiSeg and OPAL, we further run KMeans (k=10) on the first 10 PCs of the latents throughout the video and show the segmentation result at the bottom of the latent representations (Fig. 4C, D). We also show the behavioral syllable segmentation produced by Keypoint-MoSeq in Fig. 4E.
By combining the discovered motifs with real animal behavior, we assess the interpretability of each motif in the five clips (Fig. 4B). First, the right-turn behavior in clip 1 is captured by the dominance of two rightward motifs (motif 1 and motif 2). A stronger movement at the head is reflected by a higher value of motif 1. Clips 2, 3, and 4 show the mouse turning right, pausing to groom its head and ears, and then continuing to turn right. The alternate dominance of motif 3 and motif 4 aligns well with the behavior dynamics. Clip 5 shows a simultaneous behavioral mixture of moving forward and sniffing, and is captured by the equal strength of motif 5 and motif 6. Across all motifs, motif 4 appears across clips 2, 3, and 4, showing its general utility. The transitions and mixtures of behaviors are effectively reflected in the learned motifs and their temporal weights $u(t)$.
However, when we examine the segmentation and latent produced by other methods (Fig. 4C-F), we find inconsistencies. In Fig. 4C, alternate dominant behavior patterns in Clip 2-4 in the video are not reflected in its motif weights during this period, as the only dominant motif is the yellow-green one. In Fig. 4D, the motif weights are too dense to interpret. The segmentation results at the bottom of Fig. 4 C and D do not even have repeated behavior patterns and thus could not be seen as a reasonable behavior motif representation. For Keypoint-Moseq, in Fig. 4E, F, clear rightward turning in clip 1 is barely visible in syllable 3 to which it is assigned. Clips 2 and 3 are both assigned to syllable 2, even though clip 2 shows pure turning right while clip 3 is dominated by grooming movements. For clip 5, the mixture of fast moving forward and sniffing is not reflected in syllable 2.
Taken together, these results show that compared with similar approaches, MCD provides a more accurate interpretation of pose dynamics and could capture more complex behaviors through a compact, task-agnostic set of motor motifs. This offers us a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motor motifs.
Figure 4: A. AUC on training and test sets. We take an example behavior video and run our algorithm. We visualize the motif weights $u(t)$ and show the representative motifs in B. For the baseline SemiSeg, we show the latent skills and segmentation results in C. For the baseline OPAL, we show the latent skills and segmentation results in D. Then we show the segmentation results of Keypoint-MoSeq (Weinreb et al., 2024) in E and the representative motifs/syllables in F.
5 Discussion
Several limitations remain to be addressed in future work. First, the accuracy of inferred motifs is sensitive to input data quality, as occlusions or tracking errors can degrade performance. Additionally, while the framework uncovers abstract motor primitives, establishing direct correspondences between these learned "motifs" and specific neural dynamics still requires further experimental validation.
Ethics statement
Beyond advancing animal behavior research, MCD has broader implications. Positively, a better understanding of motor control mechanisms could, for instance, inform new treatments for movement disorders or inspire more adaptable AI. On the other side, extending these principles to model human behavior carries ethical risks, such as perpetuating or amplifying societal biases present in training data. A robust ethical framework is essential to mitigate such risks in the development and application of these technologies.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. All source code for model training and evaluation is included in the supplementary material, allowing independent verification and replication of our experiments. The complete set of hyperparameter values is documented in the appendix. Additionally, the data preprocessing procedures and evaluation protocols are described in the main text. These resources provide sufficient information for reproducing the results reported in this paper.
Appendix A LLM Usage
In preparing this manuscript, we employed a large language model (OpenAI ChatGPT, GPT-5) as a writing assistant. The model was used exclusively for polishing English grammar, improving clarity, and suggesting more natural phrasing in certain sections of the text. All scientific content, experimental design, analyses, and interpretations were conceived, written, and verified by the authors. The LLM was not used to generate original research ideas, analyses, or results. To ensure accuracy, all model-suggested edits were carefully reviewed and, where necessary, modified by the authors.
Appendix B Hyperparameter Setting
We train MCD using the following hyperparameters:
General hyperparameters.
  • Discount factor: $\gamma=0.99$
  • Number of epochs: $1\times 10^{6}$
  • Batch size: 256256
Motif representations.
The motif representation $\phi(s,a)\in\mathbb{R}^{d}$ and $\mu(s^{\prime})\in\mathbb{R}^{d}$ were adopted with different motif dimensions $d$ depending on the task:
  • Gridworld: $d=64$
  • Animal navigation: $d=128$
  • Animal free-moving: $d=64$
Model architectures.
  • Discrete version: $\phi$ and $\mu$ are parameterized by one-hidden-layer neural networks (hidden size =512=512).
  • Continuous version:
    • $\phi$: no hidden layer
    • $\mu$: one hidden layer (hidden size =512=512)
    • $\psi$: no hidden layer
Learning rates.
  • Discrete version:
    $\phi:1\times 10^{-3},\quad\mu:1\times 10^{-3},\quad u:3\times 10^{-4},\quad w:3\times 10^{-4}.$
  • Continuous version:
    $\psi:5\times 10^{-4},\quad\nu:5\times 10^{-4},\quad f:3\times 10^{-4},\quad u:3\times 10^{-4}.$
    During testing, $u$ and $f$ are further optimized on the new sequence using gradient descent with learning rate $1\times 10^{-3}$.
We train SemiSeg and OPAL using the following hyperparameters:
  • Discount factor: $\gamma=0.99$
  • Number of epochs: $1\times 10^{6}$
  • Batch size: $4\times 250$ (4 sequences, each of length=250 because this is an RNN-based inference model)
  • Latent dimension: $d=64$
  • Learning rate: $1\times 10^{-4}$ (tuned to get better results)
Besides, the following loss coefficients are shared across three models for interpretability results.
  • Temporal smoothness Gaussian random walk loss: 1010
  • Sparsity L1-loss: $0.1$
Appendix C Approximating Energy-Based Formulation with Low-Rank Spectral Decomposition
By simple algebra, we obtain the quadratic potential function,
$P(s^{\prime}|s,a)$ $\propto q(s^{\prime})\exp\left(\|\psi(s,a)\|^{2}/2\right)\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right).$ (9)
The term $\exp\left(-\frac{\|\psi(s,a)-\nu(s^{\prime})\|^{2}}{2}\right)$ is the Gaussian kernel, for which we apply the random Fourier feature (Dai et al., 2014; Rahimi & Recht, 2007) and obtain the spectral decomposition of Eq. 6 as
$P(s^{\prime}|s,a)=\langle\phi_{\omega}(s,a),\mu_{\omega}(s^{\prime})q(s^{\prime})\rangle_{\mathcal{N}(\omega)},$ (10)
where $\omega\sim\mathcal{N}(0,I)$ is the frequency in the Fourier domain, and
$\phi_{\omega}(s,a)$ $=\exp\left(-i\omega^{\top}\psi(s,a)\right)\exp\left(\|\psi(s,a)\|^{2}/2-\log Z(s,a)\right),$ (11)
$\mu_{\omega}(s^{\prime})$ $=\exp\left(-i\omega^{\top}\nu(s^{\prime})\right)\exp\left(\|\nu(s^{\prime})\|^{2}/2\right).$ (12)
Note that Eq. 10 needs infinite $\omega$ to calculate the expectation. To connect it to finite dimension $\phi(s,a)\in\mathbb{R}^{d},\mu(s^{\prime})\in\mathbb{R}^{d}$, we use the Monte-Carlo method to approximate it with finite samples,
$P(s^{\prime}|s,a)\approx\frac{1}{M}\sum_{i=1}^{M}\phi_{\omega_{i}}(s,a)\mu_{\omega_{i}}(s^{\prime})q(s^{\prime}).$ (13)
Introduce vectors $\phi(s,a)$ and $\mu(s^{\prime})$ such that
$\phi(s,a)$ $:=\frac{1}{\sqrt{M}}\left[\phi_{\omega_{1}}(s,a),\phi_{\omega_{2}}(s,a),...,\phi_{\omega_{M}}(s,a)\right],$ (14)
$\mu(s^{\prime})$ $:=\frac{1}{\sqrt{M}}\left[\mu_{\omega_{1}}(s^{\prime}),\mu_{\omega_{2}}(s^{\prime}),...,\mu_{\omega_{M}}(s^{\prime})\right].$ (15)
Then it’s straightforward to see that,
$\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})=\frac{1}{M}\sum_{i=1}^{M}(\phi_{\omega_{i}}(s,a)^{\top}\mu_{\omega_{i}}(s^{\prime}))\approx P(s^{\prime}|s,a).$ (16)
Hence, Eq. 6 can, in principle, yield the motif representation introduced earlier.
Appendix D Multi-task Gridworld Dataset
D.1 Dataset
To generate the dataset, we follow this procedure: 1) Use soft value iteration to compute the ground truth Q-function for each task: $Q(s,a,t)=r(s,a,t)+\log\sum_{a}\exp V(s,t)$; 2) Use the resulting Q-function to define the policy: $\pi(a|s,t)=\frac{\exp(Q(s,a,t))}{\sum_{a^{\prime}}\exp(Q(s,a^{\prime},t))}$ and sample trajectories accordingly.
Figure 5: A. State-action maps for all 64 motifs. B. Reward weight $w$ for all 64 motifs.
D.2 Learned Motifs
We visualize all original 64 motifs (Fig. 5) introduced in Sec. 4.1. It shows some meaningful patterns as mentioned before. For example, motif 0 assigns high values to those $(s,a)$ pairs leading to the middle-middle grid and the bottom-middle grid, and assigns low values to the up-left grid and the up-right grid. Thus, it is employed negatively in Task 0 (up-left reward) and Task 2 (up-right reward). However, Task 4 (middle-middle) didn’t use this motif and used motif 39 negatively instead. The complex many-to-many relationship between motifs and tasks informs us of the redundancy in the original motifs, which inspires us to use PCA to analyze the principal components of the motif space and simplify the motif weights. It could be seen from the comparison between Fig. 2D and Fig. 5A that principal components are a less redundant description of the motif space.
Appendix E Animal Navigation Behavior Dataset
E.1 Learned Motifs
In the original motifs of the labyrinth environment, multiple $(s,a)$ pairs are simultaneously activated, so it is rather hard to analyze which $(s,a)$ pairs are the most important ones that could represent the focus and function of the motif. Given the redundancy of the motif sets, as in Appendix. D, we perform PCA to analyze the principle components of the motif space and simplify the motif representation. To show the effect of PCA, we plot one motif (motif 0) before (Fig. 6A right) and after PCA (Fig. 6B right). Basically, we only want to show the most important pairs in one map and do not want low-value pairs to disturb the visualization. To determine how many $(s,a)$ pairs are important, we sort the $(s,a)$ pairs based on the value $\phi(s,a)$ in motif 0, i.e., the first feature of the output of $\phi(s,a)$ (Fig. 6A left and B left). It could be seen straightforwardly that after PCA, the motif becomes more concentrated on several $(s,a)$ pairs. We calculate the mean $\mu$ and variance $\sigma$ across all motifs and all dimensions and take $\mu+\sigma$ as the threshold, above which $(s,a)$ pairs are deemed the most important ones and are shown on the right. We show 80 pairs before PCA and 8 pairs after PCA. The number of the most important pairs in each motif is called the “effective dimension.” The effective dimension is calculated across all motifs (Fig. 6C). Paired t-test ($p=1.3\times 10^{-51}$) shows that there exists a significant decrease of effective dimensions after PCA. So the map becomes more distinct and functionally separated.
Figure 6: A. Left. The value for each $(s,a)$ pair in motif 0 before PCA. Right. The most important $(s,a)$ pairs. B. Left. The value for each $(s,a)$ pair in motif 0 after PCA. Right. The most important $(s,a)$ pairs. C. Boxplot for the effective dimensions before and after PCA.
Appendix F Animal Free-moving Behavior Dataset
F.1 Dataset and training
We split the full dataset into training and test trajectories in an 8:2 ratio. We first learn both $f$ and $u(t)$ on the training set. Then, given the learned $f$, we estimate $u(t)$ on the test set. Here, $f$ is a time-invariant model parameter shared across all time, while $u(t)$ is a time-dependent variable that must be inferred separately for each test trajectory and cannot be transferred from training.
F.2 Learned Motifs
We show all motifs learned from the 200-timestep video clip of the free-moving mouse mentioned in Sec. 4.3. We have completed the visualization of those motifs that were previously omitted due to their perceived lack of importance. Due to the increased number of displayed motifs, we had to renumber each motif. We show the present motif number above the motif motion field figure, and previous numbers (if applicable) in the parentheses following the present number.
Figure 7: A. Policy weight $u(t)$ for all 64 motifs. B. The motion field for all 64 motifs learned from the video, computed by averaging the states and actions that most strongly activate each motif.