### v3

# Learning Task-Agnostic Motifs to Capture the Continuous Nature of Animal Behavior

Abstract
Animals flexibly recombine a finite set of core motor motifs to meet diverse task demands, but existing behavior segmentation methods oversimplify this process by imposing discrete syllables under restrictive generative assumptions. To better capture the continuous structure of behavior generation, we introduce motif-based continuous dynamics (MCD) discovery, a framework that (1) uncovers interpretable motif sets as latent basis functions of behavior by leveraging representations of behavioral transition structure, and (2) models behavioral dynamics as continuously evolving mixtures of these motifs. We validate MCD on a multi-task gridworld, a labyrinth navigation task, and freely moving animal behavior. Across settings, it identifies reusable motif components, captures continuous compositional dynamics, and generates realistic trajectories beyond the capabilities of traditional discrete segmentation models. By providing a generative account of how complex animal behaviors emerge from dynamic combinations of fundamental motor motifs, our approach advances the quantitative study of natural behavior.
behavioral neuroscience, animal behavior, compositional behavior modeling, behavior segmentation, neuroethology
1 Introduction
A critical direction in animal behavior research has been identifying recurring patterns, often referred to as stereotyped behavioral syllables, like back grooming, running, and sniffing, directly from large-scale behavior recordings. Behavior segmentation methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Hsu and Yttri, 2021; Berman et al., 2014) seek to uncover such structured patterns in behavior by dividing continuous pose trajectories into discrete syllables. Classic behavior segmentation approaches can be categorized into three groups: (1) supervised classification (Marks et al., 2022; Segalin et al., 2021), (2) clustering-based methods (Hsu and Yttri, 2021; Berman et al., 2014; Whiteway et al., 2021), and (3) hidden-Markov-model(HMM) based methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Costacurta et al., 2022). The segmented syllables can then be used to build structured representations of movement for downstream neurobehavioral study.
However, existing behavior segmentation methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Hsu and Yttri, 2021; Berman et al., 2014) overlook several features of the behavior data. First, continuity, they model continuous behavior as combinations of discrete action syllables, which oversimplifies the inherently continuous nature of movement and introduces ambiguity during action transitions. Therefore, they may fail to capture the details of delicate behavior dynamics. Second, compositionality, they often extract complex coordinated body movements as abstract syllables that fail to capture how individual body parts contribute to different motions. Therefore, they fail to reveal the connections and distinctions between syllables. For example, back and side grooming both involve similar forelimb movements combined with different turning dynamics, and sniffing may occur while walking or sitting, with similar head motion but distinct lower-body patterns. Third, long-term dependency. They either ignore temporal dependency between actions or only consider a very short time window when encoding the behavior. Therefore, they often fail to capture the long-term and multi-scale property of syllables (See Appendix K for discussions on temporal dependency). Apart from these explicit features, most models are either non-generative (e.g., clustering (Hsu and Yttri, 2021; Berman et al., 2014)) or rely on restrictive generative assumptions (e.g., linear dynamics and Markov models (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022)), often leading to unrealistic synthesized behaviors.
To address these limitations, we introduce a new perspective: modeling behavior under the reinforcement learning (RL) framework. We study behavioral dynamics and motor motifs by inferring the animal’s policy through an RL-based imitation learning (IL) framework. It not only enables more realistic behavior generation through RL but also allows us to discover reusable motor motif sets to construct a policy driven by internal rewards. By viewing behavior through this lens, we gain a more flexible, generative, and interpretable understanding of motor motifs that go beyond the constraints of discrete segmentation. Note that (Aldarondo et al., 2024) also applied RL-based IL to analyze animal behavior, but without parsing long untrimmed behaviors into fine-grained, interpretable motor motifs, so their work lies outside the scope of behavior segmentation considered here.
We hypothesize that animals draw from a fixed set of core motor motifs to construct diverse movements over long behavioral trajectories (Santuz et al., 2019; Flash and Hochner, 2005). Building on this, we propose Motif-based Continuous Dynamics discovery (MCD) to parse long trajectories and uncover motifs and policies that reflect behavioral dynamics. Concretely, we learn interpretable latent representations, or motif sets, via spectral decomposition–based representation learning in RL (Dai et al., 2017; Ren et al., 2022; Shribak et al., 2024). These motifs correspond to low-level motion patterns serving as modular building blocks of behavior and can involve different body parts. For instance, face grooming (forepaw-to-face) and body grooming (torso strokes) share grooming motifs while engaging distinct body parts. These motifs can then be used to sufficiently represent policies that characterize complex high-level behaviors. Finally, we apply imitation learning to train motif-based policies from demonstrations. This framework leverages RL in two aspects: (1) motifs are inferred through RL-based representation learning, and (2) we use policies formed from motifs to characterize behavioral dynamics. As shown later, both aspects avoid any model assumptions while capturing motifs and policies that faithfully reflect behavioral dynamics.
Figure 1: A policy for a movement can be seen as a blend of “vocabularies” from a dictionary containing fundamental motor motifs.
Another key innovation in constructing policies from motifs is that each motif’s contribution evolves continuously over time, reflecting dynamic behavioral changes. Some motifs may be brief, others prolonged, and multiple motifs can be active simultaneously to build ongoing movement (Fig. 1). This flexible, compositional view cannot be achieved with traditional discrete syllables. MCD thus provides a nuanced account of how complex actions arise from dynamic motif combinations and enables testing whether fine-grained motifs depend on specific neural circuits. Compared with prior segmentation methods, MCD offers soft segmentation that captures continuous time-varying processes rather than discrete switches. Further, theoretically the horizon in RL is infinite, enabling longer temporal dependency modeling.
Note that there is a rich literature in robotics on learning skills through reinforcement learning (Appendix I). We use RL-based imitation not to compete with existing skill learning methods, as most are not suited to the behavior segmentation task for neuroscience study in this paper. Rather, our goal is to employ the RL framework as a principled way to characterize the continuous nature of animal behavior while rendering fine-grained motor motifs, analogous to how Keypoint-MoSeq relies on an SLDS framework (Weinreb et al., 2024) and VAME relies on an autoencoder framework (Luxem et al., 2022). To summarize, we contribute to behavior understanding for neuroscience and neuroethology through the following points:
  • We introduce the first RL-based IL framework for behavior segmentation, a fundamental advance since RL naturally treats behavior as a decision-making process shaped by policies and rewards. Unlike dynamics-based methods, it explains why behaviors occur, not just how they unfold.
  • Within this framework, we propose RL-based representation learning to discover motif-based policies. The learned motifs and policies do not rely on dynamics assumptions or any model assumptions. They can faithfully characterize behavioral dynamics without the mismatch issues of prior segmentation methods.
  • Our method reveals the continuous, compositional and long-term dependent nature of animal behavior, providing a nuanced understanding of how complex behaviors emerge from dynamic combinations of fundamental motor motifs. It moves beyond the discrete segmentation assumptions of existing methods.
2 Preliminaries
Markov Decision Processes (MDP). To define motifs that characterize animal behaviors, we begin by modeling the observed behavioral trajectories within the framework of MDPs. Formally, an MDP is defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,P,\rho,\gamma,H)$, where $\mathcal{S}$ denotes the state space, capturing both the environment and an animal’s condition—for example, positions of pose keypoints; $\mathcal{A}$ is the action space denoting feasible movements; $r:\mathcal{S}\times\mathcal{A}\to[0,1]$ is a reward function encoding the immediate utility toward an internal goal; $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the transition operator, with $\Delta(\mathcal{S})$ representing distributions over $\mathcal{S}$; $\rho\in\Delta(\mathcal{S})$ is the initial state distribution; $\gamma\in(0,1)$ is a discount factor; and $H$ is the time horizon. A policy $\pi:\mathcal{S}\times[H]\to\Delta(\mathcal{A})$ is a conditional distribution over actions given a state for each time-step. We assume that an animal generates pose trajectories by following such an MDP where the reward function reflects an intrinsic motivation driving behavior. The behavior is governed by a policy that seeks to maximize this internal reward.
Following standard notations, we define the value function $V(s):=\mathbb{E}\left[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the action-value function $Q(s,a)=\mathbb{E}\left[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$, which are the expected discounted cumulative rewards when executing policy $\pi$. From the above definition, we can establish the following Bellman relationship:
$Q_{h}(s,a)$ $=r(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\bigl[V_{h+1}(s^{\prime})\bigr],$
$V_{h}(s)$ $=\mathbb{E}_{a\sim\pi(\cdot\mid s)}\bigl[Q_{h}(s,a)\bigr].$ (1)
Offline Imitation Learning. We use imitation learning(IL) to find a policy $\pi$ that mimics animal behavior. In the offline IL setting, we cannot interact with the MDP environment to collect samples using policy $\pi$, but can only access a dataset of transitions sampled from the MDP by the expert, $\mathcal{D}=\left\{(s_{i},a_{i},s^{\prime}_{i})|(s,a)\sim\tau^{e},s^{\prime}\sim P(\cdot|s,a),i=1,2,...,N\right\}$, where $\tau^{e}$ is the data distribution of state-action pairs generated by the expert which is the animal in this study.
3 Motif-based Continuous Dynamics (MCD) Discovery
Given the MDP definition, we frame motif discovery from a control-theoretic perspective. In this view, motifs are the fundamental components that enable the construction of diverse policies and reward functions, and thus help explain the motivation behind observed behaviors.
Definition 3.1 (Motif Set).
Given an arbitrary transition kernel $P(\cdot|s,a)$ in an MDP, we can express it via a spectral decomposition:
$P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime}),$ (2)
where $\phi:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$, $\mu:\mathcal{S}\rightarrow\mathbb{R}^{d}$, and $q\in\Delta(\mathcal{S})$ is a parametrized probability distribution over the state space. We define the function $\phi$ as the motif set. The reward function is then parametrized linearly as $r(s,a)=\phi(s,a)^{\top}w$.
Spectral decomposition has been widely studied in RL representation learning (Ren et al., 2022; Zhang et al., 2022; Shribak et al., 2024). We adopt this approach here to define motifs given the transition kernel and define rewards accordingly. Since spectral decomposition derives latents directly from the transition kernel without model assumptions, motif learning is thus independent of model assumptions and faithfully reflects the motifs present in the behavior data.
Given the motif definition, substituting Eq. 2 and the linearized reward model into Bellman equation (Eq. 2), we get:
$Q(s,a)$ $=r(s,a)+\gamma\!\!\int\!V(s^{\prime})P(s^{\prime}|s,a)\mathrm{d}s^{\prime}$
$=\phi(s,a)^{\top}\!\!\left[w+\gamma\!\!\int\!V(s^{\prime})\mu(s^{\prime})q(s^{\prime})\mathrm{d}s^{\prime}\right]$
$=\phi(s,a)^{\top}u,\!$ (3)
where $u=w+\gamma\int V(s^{\prime})\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$. Thus, the action-value function $Q(s,a)$ can be expressed as a linear combination of motif features $\phi(s,a)$, offering a convenient way to link motifs to the policy. Following the maximum entropy reinforcement learning framework (Haarnoja et al., 2018), we assume the animal’s objective is to maximize the expected reward augmented by the policy’s entropy. Under this assumption, the optimal max-entropy policy $\pi(a|s)$ can be shown to follow:
$\pi(a|s)$ $=\arg\max_{\pi}\;\;\left[\mathbb{E}_{\pi}[Q(s,a)]+H(\pi)\right]$
$=\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\phi(s,a^{\prime})^{\top}u)},$ (4)
where the entropy $H(\pi):=-\sum_{a\in\mathcal{A}}\pi(a|s)\,\log\bigl(\pi(a|s)\bigr)$.
Proposition 3.2.
The policy in Eq. 3 is not based on any model assumption but emerges naturally as the max-entropy policy based on spectral decomposition of the transition kernel. Furthermore, the learned motifs $\phi$ can represent any max-entropy policy through an appropriate choice of $u$.
The reason we define $\phi(s,a)$ as motifs is that $\phi$ provides the linear basis for the environment transition, policies, and rewards. Policies characterize behavioral dynamics by describing action tendencies conditioned on state (how behaviors evolve). Combined with the environment dynamics $P(s^{\prime}|s,a)$, it induces the transition distribution $P(s^{\prime}|s)=\sum_{a}P(s^{\prime}|s,a)\,\pi(a|s)$, which governs the evolution of behavior trajectories, as is often directly modeled in dynamics-based methods (Wiltschko et al., 2015; Weinreb et al., 2024). Rewards, in turn, reflect the underlying driving factors of these trajectories (why behaviors evolve). Moreover, because $\phi$ is derived solely from the transition dynamics $P(s^{\prime}|s,a)$, it remains independent of any specific reward or task. In this sense, $\phi(s,a)$ encodes intrinsic, general-purpose motor motifs available to animals, while the weight vector $u$ captures task-specific modulations required to produce behavior aligned with different goals. Thus, we can interpret behavioral trajectories through the lens of motifs $\phi$.
From Def. 3.1 and Prop. 3.2, we conclude that the learned motifs and policies do not rely on model assumptions, yet the policies faithfully capture behavioral dynamics as action tendencies conditioned on state. Thus, our method is assumption-free while capturing dynamics, unlike classification/clustering methods (no dynamics) or dynamics-based methods (restrictive assumptions).
Next, we introduce how to learn $\phi(s,a)$ and $\mu(s^{\prime})$ (motif discovery), as well as $u$ (motif-based policy learning) from demonstrations. The learning procedure differs depending on the nature of the behavior data (i.e. discrete or continuous).
3.1 Discrete Version
Motif discovery. For discrete state-action spaces, there are two ways to learn the representation: (1) spectral methods such as singular value decomposition (SVD) (Golub and Reinsch, 1971; Golub and Van Loan, 2013; Trefethen and Bau, 2022); (2) spectral decomposition representation (Ren et al., 2022; HaoChen et al., 2021) with the objective $\phi(s,a),\mu(s^{\prime})=\arg\min_{\phi,\mu}||P(s^{\prime}|s,a)-\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})||^{2}$. Here we choose the second method and set the auxiliary distribution as $q(s^{\prime})=\rho(s^{\prime})=\int\tau^{e}(s,a)da$, i.e. the marginalized state distribution. The resulting motif set $\phi(s,a)$ is then used in the subsequent policy learning stage.
Motif-based policy learning. We now learn the policy $\pi(a|s)$, parameterized by Eq. 3, using maximum likelihood estimation (MLE), i.e., by optimizing the following objective to solve for $u$:
$\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\pi(a|s)\right]$
$=$ $\max_{u}\;\;\mathbb{E}_{(s,a)\sim\tau^{e}}\left[\log\frac{\exp(\phi(s,a)^{\top}u)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\phi(s,a^{\prime})^{\top}u\right)}\right].$ (5)
3.2 Continuous Version
While learning from discrete data is relatively straightforward, the continuous case presents additional challenges for two main reasons. First, in the motif discovery step (Eq. 2), the decomposition $P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ is too restrictive to capture the complexity of continuous behavioral dynamics, such as pose transitions in freely moving animals (Weinreb et al., 2024). For example, consider a common and biologically plausible behavioral model: $s^{\prime}=h(s,a)+\epsilon$, where $h$ is a dynamics function and $\epsilon$ is Gaussian noise. This additive structure, widely used in behavioral modeling, contrasts with the multiplicative form $\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, suggesting that parameterizations preserving additive relationships between $\{s,a\}$ and $s^{\prime}$ are more suitable. Second, in motif-based policy learning, for discrete datasets with a small action space, the denominator (partition function) in Eq. 3.1 is easy to compute. But for continuous data, the action space is infinite, making it infeasible to enumerate all actions and integrate. In light of these challenges, we adopt an alternative approach to learn the motifs and policy in continuous state-action spaces.
Motif discovery. We model $P(s^{\prime}|s,a)$ as an energy-based model (EBM) (Shribak et al., 2024):
$P(s^{\prime}|s,a)$ $=q(s^{\prime})\exp\left(\psi(s,a)^{\top}\nu(s^{\prime})-\log Z(s,a)\right),$
$Z(s,a)$ $=\int q(s^{\prime})\exp(\psi(s,a)^{\top}\nu(s^{\prime}))ds^{\prime},$ (6)
where $\psi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{g}$ and $\nu:\mathcal{S}\to\mathbb{R}^{g}$ are neural-network feature maps. Here, $Z(s,a)$ is an intractable partition function. Compared to the unnormalized inner-product model (Eq. 2), this EBM formulation yields smooth, normalized probabilities and stable gradients, leading to more effective and generalizable motif representations.
Proposition 3.3 (Connection to Motif Definition).
Given the EBM model in Eq. 3.2, the transition kernel can be approximated by $P(s^{\prime}|s,a)\approx\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$, where $\phi(s,a)\in\mathbb{R}^{d}$ is an explicit function of $\psi(s,a)$ and the partition function $Z(s,a)$, and $\mu(s^{\prime})\in\mathbb{R}^{d}$ is a function of $\nu(s^{\prime})$.
Appendix D contains the full proof and derivation of $\phi$ and $\mu$ in terms of $\psi$, $\nu$, and $Z$.
To learn $\psi$ and $\nu$, we employ noise-contrastive estimation (NCE) (Wu et al., 2018; Ma and Collins, 2018; Gutmann and Hyvärinen, 2010, 2012), which enables optimization of unnormalized statistical models without explicitly computing the partition function. In this way, we sidestep the intractable computation of $Z(s,a)$ in Eq. 3.2 by solving
$\max_{\psi,\,\nu}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},\\ s^{\prime}\sim P(\cdot|s,a),\\ s^{\prime\prime}_{i}\sim\rho\end{subarray}}\left[\log\frac{F(s,a,s^{\prime})}{F(s,a,s^{\prime})+\Sigma_{i=1}^{k}F(s,a,s^{\prime\prime}_{i}))}\right],$ (7)
where $F(s,a,s^{\prime})=\exp(\psi(s,a)^{\top}\nu(s^{\prime}))$, $s^{\prime}$ denotes a positive sample drawn from the transition distribution $P(\cdot|s,a)$, and $s^{\prime\prime}_{i}$ for $i=1,\ldots,k$ are negative samples from marginalized state distribution $\rho(s)=\int\tau^{e}(s,a)da$.
Connection to behavioral dynamics. With simple algebra, we obtain the quadratic potential function $P(s^{\prime}|s,a)\propto q(s^{\prime})\cdot\exp\left(\|\psi(s,a)\|^{2}/2\right)\cdot\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right)\cdot\left(\|\nu(s^{\prime})\|^{2}/2\right)$ from Eq. 3.2. By enforcing unit-norm constraints $\|\psi(s,a)\|^{2}=\|\nu(s^{\prime})\|^{2}=1$, assuming $Z(s,a)$ as a constant and $q(s^{\prime})$ as uniform distribution, as well as taking $\nu$ to be the identity map, we obtain a generalized Gaussian form: $P(s^{\prime}|s,a)=\frac{1}{Z}\exp\left(-\|s^{\prime}-\psi(s,a)\|^{2}\right)$, which aligns with commonly adopted assumptions in animal behavior modeling discussed earlier. Thus, Eq. 3.2 offers a more general framework that extends traditional dynamics models for studying behavior. Moreover, in practice, using Eq. 3.2 results in a unimodal distribution over $s^{\prime}$, which closely matches the empirical structure observed in $P(s^{\prime}|s,a)$ from animal behavior data. Thus, while Eq. 2 is theoretically valid in continuous domains, we adopt the parameterization in Eq. 3.2 for continuous state and action spaces, as it more effectively supports the learning of motif representations underlying animal behavior.
Motif-based policy learning. After we obtain the representation $\psi(s,a)$ and $\nu(s^{\prime})$ from Eq. 7, we could theoretically get motif sets $\phi(s,a)$ expressed as the function of $\psi(s,a)$ and a normalizing term $Z(s,a)$ (see Appendix D). However, since in practice $Z(s,a)$ remains intractable, even with the optimal $\psi$ it is still hard to obtain $\phi$ exactly. Therefore, we introduce a mapping $f:\psi\rightarrow\phi$, parameterized with a neural network, and learn it via policy learning. Our aim is to learn a function $f$ so that $\phi=f(\psi)$ yields optimal basis functions of policy that best account for the animal behavior data. By applying $\phi=f(\psi)$ to Eq. 3, we obtain $Q(s,a)=f\bigl(\psi(s,a)\bigr)^{\top}u$.
As mentioned earlier, learning both $f$ and $u$ using the MLE objective in Eq. 3.1 becomes intractable for continuous data, as the denominator involves integration over an unbounded continuous action space. This brings us back to the challenge of estimating an unnormalized energy function, $\pi(a|s)\propto\exp(Q(s,a))\propto\exp(f(\psi(s,a))^{\top}u)$. Thus, it’s reasonable to apply NCE here again,
$\max_{f,\,u}\;\;\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\tau^{e},\\ (s_{i}^{\prime},a^{\prime}_{i})\sim\tau^{e}\end{subarray}}\left[\log\frac{G(s,a)}{G(s,a)+\Sigma_{i=1}^{k}G(s,a^{\prime}_{i})}\right],$ (8)
where $G(s,a)=\exp(f(\psi(s,a))^{\top}u)$, $\{s,a\}$ are positive samples and $\{s^{\prime}_{i},a^{\prime}_{i}\}$ are negative samples. Identifiability discussions are provided in Appendix J.
3.3 Understanding Animal Behavior via Motifs
In this section, we discuss how motifs $\phi$ and weights $u$ can be used to describe animal behavior trajectories, by allowing $u$ to vary across tasks or time points and dynamically modulate each motif. We consider two scenarios: (1) discrete state-action spaces in a multi-task setting, and (2) continuous state-action spaces in a time-varying setting. In either case, $u(t)$ would be a matrix where each column corresponds to one weight for one task $t$ or time point $t$.
Scenario (1): Consider a mouse navigating a maze, where trajectories are discretized into a finite state and action space. We assume the animal switches between $T$ strategies, each associated with a unique reward function. We first learn shared motifs $\phi(s,a)$ using Eq. 2, then fit task-specific policies $\pi(a|s,t)$ using Eq. 3.1, where each task $t$ is one of the $T$ strategies. This yields $T$ sets of weights $u(t)$, one per task, which share a common motif set $\phi(s,a)$ across tasks.
Scenario (2): Consider a mouse freely moving in the cage, where the states are the continuous coordinates of the body parts, and actions are the continuous velocities. We learn $\psi(s,a)$ from pose trajectories using Eq. 7, and learn $u(t)$ and $f$ using Eq. 8. Finally, we get motor motifs in $\phi(s,a)=f(\psi(s,a))$, and how they can construct time-varying policy $\pi(a|s,t)$ through weights $u(t)$. The policy is finally represented by a generalized linear combination of basis motifs, producing two advantages: (1) Continuity. Unlike models with abrupt discrete switches, our approach features continuous-time formulation with smoothly evolving motifs, and can capture subtle behavioral changes more faithfully; (2) Compositionality. Unlike models assigning one discrete label to one timestep, our approach provides a compositional view and allow two motifs to be active simultaneously. For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side. (3) Long-term dependency. Unlike models only considering short time windows, our approach has multi-scale temporal dependency and infinite horizon in theory (Appendix K).
Taken together, MCD offers a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a continuous, compositional and multi-scale time-varying description of animal trajectories. In Experiment (Sec. 4), we would show (1) what motifs we have learned, and (2) how they are used to construct the final policy, on one simulation datasets and two real animal behavior datasets.
3.4 Reward Recovery
After estimating $u(t)$ as the motif weights for policy construction, we can further infer $w(t)$ for reward representation $r(s,a,t)=\phi(s,a)^{\top}w(t)$ as $w(t)=u(t)-\gamma\int V(s^{\prime},t)\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$, where $V(s,t)=\log\sum_{a}\exp Q(s,a,t)$. This allows us to recover the time-varying reward function $r(s,a,t)$ used by animals. Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) (Ziebart et al., 2008). In the context of animal behavior (Ke et al., 2025; Zhu et al., 2023; Ashwood et al., 2022), identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal decision-making processes. Since $V(s,t)$ can only be easily computed in closed form in discrete and finite settings, we validate our method by visualizing the inferred rewards in the first two experiments.
4 Experiments
4.1 Application to simulated multi-task gridworld
Figure 2: A. State-action map: each of the nine grids is divided into four cells representing action values (Up, Down, Left, Right). In task $i$, high reward is assigned to $(s,a)$ pairs moving toward the $i$th location. B. Explained variance of the top 15 principal components; variance drops near zero after PC7. C. Left: true rewards for all 9 tasks (yellow = high, green = low). Right: recovered rewards. D. Reward weight $w$ and top 8 PC motifs from the $\phi$ matrix. Reward weights indicate the contribution of the top 8 PC motifs to each task. In the PC motif plot, red indicates positive feature values, blue indicates negative.
The gridworld consists of a $3\times 3$ lattice, and each of the nine states allows four possible actions: Up, Down, Left, and Right (Fig. 2A). In task $t\in\{0,...,8\}$, a fixed reward is assigned to the $(s,a)$ pairs that move toward state $t$. Fig. 2C (left) shows the ground truth of reward functions for all nine tasks. We collect a dataset where in each episode, the agent starts from a random state and navigates to task-specific state $t$. See Appendix E for more details for data generation.
As described in Sec. 3.1, we learn a set of latent motifs and use them to construct the task-specific policy $\pi(a|s,t)$. Because the computational complexity only scales linearly with the number of motifs, to cover the motif space as much as possible, we select a large number for motif dimension $d=64$. (See Appendix L.1 for discussions of motif dimensions.) Visualizations of these motifs are shown in the Appendix E. Using the learned motifs, we recover the policy and further infer the reward function, as described in Sec. 3.4, with the form: $r(s,a,t)=\phi(s,a)^{\top}w(t)$. This recovered reward (Fig. 2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned motifs are sufficient for accurately reconstructing the reward function from behavior data.
To better interpret the learned motifs and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig. 2B). These components are basis vectors spanning the motif space. To check if they exhibit interpretable patterns, we visualize the top 8 PC motifs and their corresponding task-specific coefficients $w(t)$ in Fig. 2D. For instance, in motif 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values. This motif corresponds to moving away from the middle-right grid toward the bottom-left. Accordingly, we observe at Column 0 in $w$ a high value at Row 6 and a low value at Row 5, suggesting that Task 6 requires this motif to construct its reward function. Similar interpretations can be made for other motifs and tasks. Therefore, the learned motifs are effectively deployed to recover reward functions.
4.2 Application to animal navigation behavior
Dataset and model setup.
We next evaluate our method on a real animal behavior dataset from (Rosenberg et al., 2021). In this experiment, a thirsty mouse is trained to start from the home cage and navigate to the water port in a binary-tree maze (Fig. 3A). The state space is defined by the mouse’s location on the tree. The actions include moving to its left parent, right parent, left child, and right child. Although the mouse’s behavior is primarily driven by water foraging, it also exhibits exploration of unvisited areas and returns to the home cage for shelter, which cannot be simply captured by a single reward function. Studying this dataset allows us to discover reusable and interpretable motifs shared across multiple competing motivations.
We first apply the segmentation algorithm from (Ke et al., 2025) to divide long behavior trajectories into three interpretable tasks: water seeking, home seeking, and exploration, and infer their respective reward functions (Fig. 3A, top row). While the mouse’s true internal reward functions remain unknown, we treat these inferred rewards as effective ground truth since they generate policies that closely replicate the observed behavior. Then we apply our model to identify motifs shared across these three reward functions (discrete version, Sec. 3.1). Again, we choose the motif dimension $d=127$ to completely cover the motif space. See Appendix L.1 for more details on the selection of $d$.
Figure 3: A. True and recovered rewards for the three tasks. B. Two dominant motifs for the water and home tasks respectively. Each motif indicates that taking a specific action (orange arrow) toward the dark purple state yields a high value. C. $w$ values for all tasks, colored by task; blue stars highlight high-weight motifs for water seeking, and orange for home seeking, all shown in (B).
Results.
To assess model performance, we estimate the task-specific weights $w(t)=u(t)-\gamma\int V(s^{\prime},t)\mu(s^{\prime})q(s^{\prime})ds^{\prime}$ (Eq. 3) and reconstruct the reward function as $r(s,a,t)=\phi(s,a)^{\top}w(t)$. The recovered reward functions align closely with the ground truth (Fig. 3A, bottom row). In the water-seeking task, the recovered reward has peaks near the water port and along the path leading to it. In home-seeking, a distinct peak appears at the home cage. In exploring, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
We further visualize the learned motifs by applying PCA to obtain PC motifs. See Appendix F for all raw motifs. Fig. 3B displays two top-contributing motifs for the water and home task, selected based on the peak of linear weights $w$ (Fig. 3C). The water motifs promote movement towards the water port, while the home motifs promote movement towards the cage. Moreover, most motifs are dominated by one task rather than shared across them (Fig. 3C), reflecting competitive motivations.
Unlike previous reward discovery methods on this dataset that lack motifs  (Ashwood et al., 2022; Ke et al., 2025), these results show that our model can learn one single set of interpretable motifs to recover multiple reward functions, revealing how local decision processes combine to produce complex strategies.
4.3 Application to animal free-moving behavior
Figure 4: A. Human annotated label prediction accuracy and AUC on test sets. AUC experiments are repeated for 10 sequences with 100 timesteps for each sequence. The negative samples are sampled 1000 times for each sequence. We take an example behavior video and run our algorithm. We visualize the motif weights $u(t)$ and show the representative motifs in B. For the baseline SemiSeg, we show the latent skills and segmentation results in C. For the baseline OPAL, we show the latent skills and segmentation results in D. Then we show the segmentation results of Keypoint-MoSeq (Weinreb et al., 2024) in E and the representative motifs/syllables in F.
Dataset and model setup.
To verify MCD’s generalizability, we apply our method to a continuous dataset of free-moving mouse behaviors (Weinreb et al., 2024). The state corresponds to the $x,y$ coordinate of eight body parts, including the head, nose, both ears, and four spine nodes. The action is defined as the body part velocity, $a_{t}=(s_{t+1}-s_{t})/\delta t$. We set the time interval to $\delta t=1$ ($1/30s$ in the real world). Studying this dataset allows us to ask whether free-moving behaviors, which consist of mixtures of grooming, locomotion, and postural adjustments, can be combined through a compact set of motifs.
For this continuous dataset, we use NCE to learn the motifs $\phi(s,a)$ (Sec. 3.2) and the time-dependent weights $u(t)$ (Sec. 3.3). We place a Gaussian random walk prior on $u(t)$ to ensure temporal smoothness: $u(t)\sim\mathcal{N}(u(t-1),\sigma^{2}I)$. We optimize $f$ and $u(t)$ using coordinate descent to improve stability. Since here our focus is on understanding the learned motifs and animal dynamics, we do not perform IRL but focus on interpreting $\phi$ and $u(t)$. We find in practice that the performance grows slowly once the motif dimension exceeds $d=64$ (Appendix L.1), so we choose this in practice. See Appendix G for more training details.
We compare our MCD method with two representative behavior segmentation approaches: (1) Keypoint-MoSeq (Weinreb et al., 2024), an HMM-based method; and (2) SemiSeg (Whiteway et al., 2021), a clustering-based method. Note that we only select one HMM-based method, since Keypoint-MoSeq methodologically extends the autoregressive hidden Markov model (AR-HMM, as in MoSeq (Wiltschko et al., 2015)), and is the state-of-the-art approach, outperforming a series of work (B-SOiD (Hsu and Yttri, 2021), VAME (Luxem et al., 2022), and MotionMapper (Berman et al., 2014)). We also include OPAL (Ajay et al., 2021), an autoencoder-based motif learning algorithm from robotics, as a baseline to show our advantages, and that robotics approaches are ill-suited for behavioral segmentation in neuroscience.
Results.
To evaluate whether the inferred latents $u(t)$ are meaningful, we fit a two-layer neural network to decode the human-annotated labels. MCD achieves the highest test accuracy (Fig. 4A top), suggesting motifs recovered by MCD align better with actual animal behavior dynamics. More details are provided in Appendix H.
To evaluate the model’s ability to discriminate real trajectories from mismatch trajectories, we calculate the area under the receiver operating characteristic curve (AUC), which allows direct comparison between our unnormalized energy function and other models’ normalized likelihood. Given $(s,a),(s^{\prime},a^{\prime})\sim\tau^{e}$, we define positive samples as $(s,a)$ and negative samples as mismatch pairs $(s,a^{\prime})$. For the choice of the prediction score, we use (1) action log-likelihood for Keypoint-MoSeq; (2) negative energy function $\phi(s,a)^{\top}u(t)$ for MCD; (3) Gaussian log-likelihood ($\sigma=1$) for SemiSeg; and (4) action log-likelihood for OPAL. Our model achieves the highest AUC (paired t-test, $p<0.05$ for every baseline, Fig. 4A bottom), demonstrating the strongest ability to distinguish real trajectories and accurately capture continuous dynamics through compact motifs.
Beyond quantitative comparisons, we also qualitatively visualize and interpret $\phi(s,a)$ associated with example pose dynamics. For MCD, we examine the time-varying $u(t)$ of a long animal behavior video (length=250), from which we extract five clips (length=5), each of which corresponds to a unique mixture of motifs (Fig. 4B). For each clip, we show the top 1-2 most dominant motifs, and display them using average pose and actions, computed by averaging the states and actions that most strongly activate that motif. Each motif is labeled with its meaning. A more comprehensive visualization is included in Appendix G. On the same video, we show the latents inferred by SemiSeg (Fig. 4C) and OPAL (Fig. 4D). For clarity, we further run KMeans (k=10) on the first 10 PCs of the latents and show the clustering result at the bottom. The segmentation produced by Keypoint-MoSeq is shown in Fig. 4E.
By combining the discovered motifs with real animal behavior, we assess the interpretability of each motif in the five clips (Fig. 4B). First, the right-turn behavior in clip 1 is captured by the dominance of two rightward motifs (motif 1 and motif 2). A stronger movement at the head is reflected by a higher value of motif 1. Clips 2, 3, and 4 show the mouse turning right, pausing to groom its head and ears, and then continuing to turn right. The alternate dominance of motif 3 and motif 4 aligns well with the behavior dynamics. Clip 5 shows a simultaneous behavioral mixture of moving forward and sniffing, and is captured by the equal strength of motif 5 and motif 6. Across all motifs, motif 4 appears across clips 2, 3, and 4, showing its general utility. The transitions and mixtures of behaviors are effectively reflected in the learned motifs and their temporal weights $u(t)$.
However, other models show inconsistencies. In SemiSeg (Fig. 4C), alternate dominant behavior patterns in Clip 2-4 in the video are not reflected in its motif weights during this period, as the only dominant motif is the yellow-green one. OPAL (Fig. 4D) lacks dominant and specific motifs for interpretation. Moreover, these two models do not even have segmentation label that appears again, suggesting their latents can not decompose behaviors to a simple subset. For Keypoint-MoSeq, in Fig. 4E, F, clear rightward turning in clip 1 is barely visible in syllable 3 to which it is assigned. Clips 2 and 3 are both assigned to syllable 2, even though clip 2 shows pure turning right while clip 3 is dominated by grooming movements. For clip 5, the mixture of fast moving forward and sniffing is not reflected in syllable 2.
Taken together, these results show that compared with baselines, MCD could accurately capture complex behaviors through a compact, task-agnostic set of motor motifs, offering a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motifs. We provide ablation studies in Appendix C to show that MCD is robust to motif dimension $d$, policy noise distribution, dataset partition, and Gaussian random walk prior. We also evaluate the generative quality of MCD by performing several rollouts in Appendix 16.
5 Discussion
We talk about the relations of our work to neuroscience and ethology in Appendix M. Several limitations remain to be addressed in future work. First, the accuracy of inferred motifs is sensitive to input data quality, and tracking errors can degrade performance. Additionally, while the framework uncovers abstract motor primitives, establishing direct correspondences between these learned ”motifs” and specific neural dynamics still requires further experimental validation.
Impact statement
Beyond advancing animal behavior research, MCD has broader implications. Positively, a better understanding of motor control mechanisms could, for instance, inform new treatments for movement disorders or inspire more adaptable AI. On the other side, extending these principles to model human behavior carries ethical risks, such as perpetuating or amplifying societal biases present in training data. A robust ethical framework is essential to mitigate such risks in the development and application of these technologies.
Appendix A Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. All source code for model training and evaluation is included in the supplementary material, allowing independent verification and replication of our experiments. The complete set of hyperparameter values is documented in the appendix. Additionally, the data preprocessing procedures and evaluation protocols are described in the main text. These resources provide sufficient information for reproducing the results reported in this paper.
Appendix B LLM Usage
In preparing this manuscript, we employed a large language model (OpenAI ChatGPT, GPT-5) as a writing assistant. The model was used exclusively for polishing English grammar, improving clarity, and suggesting more natural phrasing in certain sections of the text. All scientific content, experimental design, analyses, and interpretations were conceived, written, and verified by the authors. The LLM was not used to generate original research ideas, analyses, or results. To ensure accuracy, all model-suggested edits were carefully reviewed and, where necessary, modified by the authors.
Appendix C Hyperparameter Setting
We train MCD using the following hyperparameters:
General hyperparameters.
  • Discount factor: $\gamma=0.99$
  • Number of epochs: $1\times 10^{6}$
  • Batch size: 256256
Motif representations.
The motif representation $\phi(s,a)\in\mathbb{R}^{d}$ and $\mu(s^{\prime})\in\mathbb{R}^{d}$ were adopted with different motif dimensions $d$ depending on the task:
  • Gridworld: $d=64$
  • Animal navigation: $d=127$
  • Animal free-moving: $d=64$
Model architectures.
  • Discrete version: $\phi$ and $\mu$ are parameterized by one-hidden-layer neural networks (hidden size =512=512).
  • Continuous version:
    • $f$: no hidden layer
    • $\nu$: one hidden layer (hidden size =512=512), with one normalization layer to ensure the L2-Norm of the output is 1.
    • $\psi$: no hidden layer, with one normalization layer to ensure the L2-Norm of the output is 1.
  • In both cases, $u$ and $w$ are just two matrices, with each column corresponding to $u(t)$ or $w(t)$ at a specific task or timepoint.
  • Activation function: For all networks, if it has at least one hidden layer, then all of its activation functions are Exponential Linear Unit (ELU).
Learning rates.
  • Discrete version:
    $\phi:1\times 10^{-3},\mu:1\times 10^{-3},u:3\times 10^{-4},w:3\times 10^{-4}.$
  • Continuous version:
    $\psi:5\times 10^{-4},\nu:5\times 10^{-4},f:3\times 10^{-4},u:3\times 10^{-4}.$
    During testing, $u$ and $f$ are further optimized on the new sequence using gradient descent with learning rate $1\times 10^{-3}$.
We train SemiSeg and OPAL using the following hyperparameters:
  • Discount factor: $\gamma=0.99$
  • Number of epochs: $1\times 10^{6}$
  • Batch size: $4\times 250$ (4 sequences, each of length=250 because this is an RNN-based inference model)
  • Latent dimension: $d=64$
  • Learning rate: $1\times 10^{-4}$ (tuned to get better results)
Besides, the following loss coefficients are shared across three models for interpretability results.
  • Temporal smoothness Gaussian random walk loss: 1010
  • Sparsity L1-loss: $0.1$
Appendix D Approximating Energy-Based Formulation with Low-Rank Spectral Decomposition
In this part, we show in detail the connection between the EBM formulation (Eq. 3.2) and the low-rank spectral decomposition formulation (Eq. 2) of the transition kernel $P(s^{\prime}|s,a)$.
From Eq. 3.2, by simple algebra, we obtain the quadratic potential function,
$P(s^{\prime}|s,a)\propto$ $q(s^{\prime})\cdot\exp\left(\|\psi(s,a)\|^{2}/2\right)\cdot\exp\left(\|\nu(s^{\prime})\|^{2}/2\right)$
$\cdot\exp\left(-\|\psi(s,a)-\nu(s^{\prime})\|^{2}/2\right).$ (9)
The term $\exp\left(-\frac{\|\psi(s,a)-\nu(s^{\prime})\|^{2}}{2}\right)$ is the Gaussian kernel, for which we apply the random Fourier feature (Dai et al., 2017; Rahimi and Recht, 2007) and obtain the spectral decomposition of Eq. 3.2 as
$P(s^{\prime}|s,a)=\langle\phi_{\omega}(s,a),\mu_{\omega}(s^{\prime})q(s^{\prime})\rangle_{\mathcal{N}(\omega)},$ (10)
where $\omega\sim\mathcal{N}(0,I)$ is the frequency in the Fourier domain, and
$\phi_{\omega}(s,a)=$ $\exp\left(-i\omega^{\top}\psi(s,a)\right)$
$\cdot\exp\left(\|\psi(s,a)\|^{2}/2-\log Z(s,a)\right),$ (11)
$\mu_{\omega}(s^{\prime})=$ $\exp\left(-i\omega^{\top}\nu(s^{\prime})\right)\cdot\exp\left(\|\nu(s^{\prime})\|^{2}/2\right).$ (12)
Note that Eq. 10 needs infinite $\omega$ to calculate the expectation. To connect it to finite dimension $\phi(s,a)\in\mathbb{R}^{d},\mu(s^{\prime})\in\mathbb{R}^{d}$, we use the Monte-Carlo method to approximate it with finite samples,
$P(s^{\prime}|s,a)\approx\frac{1}{M}\sum_{i=1}^{M}\phi_{\omega_{i}}(s,a)\mu_{\omega_{i}}(s^{\prime})q(s^{\prime}).$ (13)
Introduce vectors $\phi(s,a)$ and $\mu(s^{\prime})$ such that
$\phi(s,a)$ $:=\frac{1}{\sqrt{M}}\left[\phi_{\omega_{1}}(s,a),\phi_{\omega_{2}}(s,a),...,\phi_{\omega_{M}}(s,a)\right],$ (14)
$\mu(s^{\prime})$ $:=\frac{1}{\sqrt{M}}\left[\mu_{\omega_{1}}(s^{\prime}),\mu_{\omega_{2}}(s^{\prime}),...,\mu_{\omega_{M}}(s^{\prime})\right].$ (15)
Then it’s straightforward to see that,
$\phi(s,a)^{\top}\mu(s^{\prime})q(s^{\prime})$ $=\frac{1}{M}\sum_{i=1}^{M}(\phi_{\omega_{i}}(s,a)^{\top}\mu_{\omega_{i}}(s^{\prime}))$
$\approx P(s^{\prime}|s,a).$ (16)
Hence, Eq. 3.2 can, in principle, yield the motif representation introduced earlier.
Appendix E Multi-task Gridworld Dataset
E.1 Dataset
To generate the dataset, we follow this procedure: 1) Use soft value iteration to compute the ground truth Q-function for each task: $Q(s,a,t)=r(s,a,t)+\log\sum_{a}\exp V(s,t)$; 2) Use the resulting Q-function to define the policy: $\pi(a|s,t)={\exp(Q(s,a,t))}/{\sum_{a^{\prime}}\exp(Q(s,a^{\prime},t))}$ and sample trajectories accordingly.
Figure 5: A. State-action maps for all 64 motifs. B. Reward weight $w$ for all 64 motifs.
E.2 Learned Motifs
We visualize all original 64 motifs (Fig. 5) introduced in Sec. 4.1. It shows some meaningful patterns as mentioned before. For example, motif 0 assigns high values to those $(s,a)$ pairs leading to the middle-middle grid and the bottom-middle grid, and assigns low values to the up-left grid and the up-right grid. Thus, it is employed negatively in Task 0 (up-left reward) and Task 2 (up-right reward). However, Task 4 (middle-middle) didn’t use this motif and used motif 39 negatively instead. The complex many-to-many relationship between motifs and tasks informs us of the redundancy in the original motifs, which inspires us to use PCA to analyze the principal components of the motif space and simplify the motif weights. It could be seen from the comparison between Fig. 2D and Fig. 5A that principal components are a less redundant description of the motif space.
Appendix F Animal Navigation Behavior Dataset
F.1 Learned Motifs
In the original motifs of the labyrinth environment, multiple $(s,a)$ pairs are simultaneously activated, so it is rather hard to analyze which $(s,a)$ pairs are the most important ones that could represent the focus and function of the motif. Given the redundancy of the motif sets, as in Appendix E, we perform PCA to analyze the principle components of the motif space and simplify the motif representation. To show the effect of PCA, we plot one motif (motif 0) before (Fig. 6A right) and after PCA (Fig. 6B right). Basically, we only want to show the most important pairs in one map and do not want low-value pairs to disturb the visualization. To determine how many $(s,a)$ pairs are important, we sort the $(s,a)$ pairs based on the value $\phi(s,a)$ in motif 0, i.e., the first feature of the output of $\phi(s,a)$ (Fig. 6A left and B left). It could be seen straightforwardly that after PCA, the motif becomes more concentrated on several $(s,a)$ pairs. We calculate the mean $\mu$ and variance $\sigma$ across all motifs and all dimensions and take $\mu+\sigma$ as the threshold, above which $(s,a)$ pairs are deemed the most important ones and are shown on the right. We show 80 pairs before PCA and 8 pairs after PCA. The number of the most important pairs in each motif is called the “effective dimension.” The effective dimension is calculated across all motifs (Fig. 6C). Paired t-test ($p=1.3\times 10^{-51}$) shows that there exists a significant decrease of effective dimensions after PCA. So the map becomes more distinct and functionally separated.
Figure 6: A. Left. The value for each $(s,a)$ pair in motif 0 before PCA. Right. The most important $(s,a)$ pairs. B. Left. The value for each $(s,a)$ pair in motif 0 after PCA. Right. The most important $(s,a)$ pairs. C. Boxplot for the effective dimensions before and after PCA.
Appendix G Animal Free-moving Behavior Dataset
G.1 Dataset and training
We split the full dataset into training and test trajectories in an 8:2 ratio. We first learn both $f$ and $u(t)$ on the training set. Then, given the learned $f$, we estimate $u(t)$ on the test set. Here, $f$ is a time-invariant model parameter shared across all time, while $u(t)$ is a time-dependent variable that must be inferred separately for each test trajectory and cannot be transferred from training.
G.2 Learned Motifs
We show all motifs learned from the 200-timestep video clip of the free-moving mouse mentioned in Sec. 4.3. We have completed the visualization of those motifs that were previously omitted due to their perceived lack of importance. Due to the increased number of displayed motifs, we had to renumber each motif. We show the present motif number above the motif motion field figure, and previous numbers (if applicable) in the parentheses following the present number.
Figure 7: A. Policy weight $u(t)$ for all 64 motifs. B. The motion field for all 64 motifs learned from the video, computed by averaging the states and actions that most strongly activate each motif.
Appendix H Evaluation details on a Human-Annotated Dataset.
We verify our method on a supervised behavior label dataset. Due to the substantial cost and labor demands of manual annotation, large-scale, high-quality animal behavior datasets with human labels remain limited. Commercial annotation tools (e.g., HomeCageScan) are also expensive, further constraining availability. To support quantitative evaluation, we therefore included a human-annotated subset from the freely moving animal behavior dataset we mentioned last section. Specifically, we selected 200 short video clips (10 recording sessions; in each session we randomly sampled 20 clips, with 200 frames each). Each clip was manually annotated with behavior segments for each frame, resulting in 200 per-frame labels across six categories: Walking, Sniffing/Grooming, Turning left, Turning right, Rearing, and Resting. Manual annotation required approximately 40 hours of expert effort. The dataset is split into a training set and a test set with a ratio $8:2$.
As before, we evaluated four models on this dataset: MCD, Keypoint-MoSeq, SemiSeg, and OPAL. Because the latent variables or syllables inferred by these models do not necessarily align with human-defined labels, we trained an additional decoder to map model-specific latent representations to the ground-truth categories. For each model, the decoder input was the motif weights $u(t)$ (MCD), the inferred one-hot syllable (Keypoint-MoSeq), or the latent embedding $z$ (SemiSeg and OPAL). The decoder was implemented as a two-layer neural network with ReLU activation, optimized using the Adam optimizer and cross-entropy loss (learning rate =0.001=0.001, hidden size=100=100), following standard settings in the scientific computation package scikit-learn. The resulting classification accuracies on the held-out test set (Fig. 4A) demonstrate that MCD substantially outperforms all baselines, achieving near-perfect accuracy. This indicates that the motifs recovered by MCD align better with actual animal behavior dynamics.
Appendix I Connections to Relevant Algorithms in RL
Motor primitives learning. Animal behavior research has focused on identifying motor primitives directly from large-scale recordings. Supervised classification (Marks et al., 2022; Segalin et al., 2021), clustering-based analysis (Hsu and Yttri, 2021; Berman et al., 2014), and HMM-based methods (Wiltschko et al., 2015; Weinreb et al., 2024; Luxem et al., 2022; Whiteway et al., 2021) label behavior into discrete syllables and analyze neural correlates of transitions. While these approaches reveal stereotyped motifs such as walking or grooming, they typically treat primitives as rigid and exclusive, overlooking the continuous, compositional structure of behavior and its adaptation across contexts.
In robotics, movement primitives have long been studied (Paraschos et al., 2013; Saveriano et al., 2023; Lioutikov et al., 2017). ProMP (Paraschos et al., 2013) extracts primitives from demonstrations but requires labeled skills/motifs, while ProbS (Lioutikov et al., 2017) jointly infers behavioral segmentation and a primitive library via Expectation-Minimization (EM), enabling unsupervised skill discovery. These approaches resemble dynamics-based behavior segmentation in animal behavior research.
All methods mentioned here lack an RL perspective that links primitives to policies and reward-driven behavior. Our framework instead leverages offline RL-based imitation learning to uncover motif-based policies that capture both modular primitives and their sequential composition, and provides a principled way to explain why behaviors occur, not just how they unfold, and how reusable motifs/skills contribute to them in a generative decision-making process.
Offline Imitation learning. Offline imitation learning presents the problem of learning a policy from fixed demonstrations when access to environments is impossible. Simple behavior cloning can be performed offline, but fails to generalize well in some cases because it does not consider dynamics or environment structure limits. ValueDICE (Kostrikov et al., 2019) considers the dynamics of the training data and learns a policy that minimizes the KL-divergence between the state-action occupancies generated by the policy and of the original dataset. But the adversarial optimization of policy and Q-functions introduces instability in training. IQL (Garg et al., 2021) avoids adversarial training by directly parameterizing the policy in terms of the $Q$-function, $\pi(a|s)=\exp(Q(s,a))/Z$. However, all of these algorithms learn policies as unstructured functions. As the animal behavior structure is highly modular and stereotyped, it is more appropriate to employ a hierarchical motif-based policy to model the data.
Motif/Skill discovery. Unsupervised skill discovery (DIAYN (Eysenbach et al., 2019), BeCL (Yang et al., 2023), DADS (Sharma et al., 2020), InfoGAIL (Li et al., 2017), Directed-InfoGAIL (Sharma et al., 2018), SkillBlender (Kuang et al., 2025), ASE (Peng et al., 2022)) to find a high-level abstraction of actions has been an effective strategy for online RL and imitation learning. However, they are limited to online settings when the skills can only be refined through interacting with the environment, which restricts their application to large-scale offline datasets. Recent offline skill discovery methods include OPAL (Ajay et al., 2021), SPiRL (Pertsch et al., 2021a), and SkiLD (Pertsch et al., 2021b) which use an autoencoder to encode trajectories into latent skills $z$; and PARROT (Singh et al., 2021) which uses a flow-based model to learn a behavior prior. In these models, the latent $z$ is later used to generate the policy $\pi(a|s,z)$ in a non-linear way. Our paper, instead, employs a generalized-linear structure of the policy $\pi(a|s,z)\propto\exp(\phi(s,a)^{\top}u(z=t))$ which provides better interpretability than the policy network. This interpretability is essential for addressing downstream neuroscience questions.
Linear structure of environment/policy. Another line of work (HILP (Park et al., 2024), FB (Touati and Ollivier, 2021), USFA (Borsa et al., 2019), RaMP (Chen et al., 2023)) based on successor features (SF) also uses a generalized linear structure to model the motif-based policy as $\pi(a|s,z)\propto\exp(F(s,a,z)^{\top}z)$ where $F(s,a,z)$ is SF under a certain motif $z$. The main concern is that they cannot separate motif representation $z$ from state-action representation in SF, while our work can (Eq. 3). Therefore, their motifs depend on the specific task or timepoint, while our motifs/skills are general representations shared across tasks and timepoints. The idea that task-agnostic motifs are combined adaptively to form new policies aligns better with the need of interpretability and scientific discovery: we would like to look for neural signals responsible for relatively fixed time-agnostic behavior patterns, to help us better understand the animal behavior.
TRAIL (Yang et al., 2022) adopts a linear decomposable environment as MCD when learning the latent skills. However, the latent-conditioned policy $\pi(a|s,z)$ is not linear, but parametrized by a neural network. SkillBlender (Kuang et al., 2025) adopts a linear decomposable policy, and uses linear combinations of lower-level controller outputs as the final actions $a_{t}=\sum_{i}k_{i}a^{i}_{t}$, while we use the linear combinations of different motifs to generate the final state-value functions $Q_{t}(s,a)=\sum_{i}\phi_{i}(s,a)u^{i}_{t}$ rather than direct actions. We believe a naive mixture of the low-level controller is less biologically realistic in modeling animal behavior than mixtures of state-value functions. The latter can find supporting evidence in neuroscientific literature (Makino, 2023). Besides, their transition kernel $P(s^{\prime}|s,a)$ is not linear.
Compared to them, our work assumes generalized linear structures for both the policy and the environment, providing better interpretability. This shared motif is more fundamental in revealing the basic structure of the animal’s intentions.
Our contributions. Most existing skill-learning methods do not apply to the behavior understanding scenario considered here. We study offline data without supervision or task annotations—no explicit goals, labels, or trajectory segmentation—only long, unstructured behavior sequences. The challenge is threefold: (1) to discover the basic motor skills, (2) to determine how these skills compose spatially and temporally within long trajectories in an offline setting, and (3) to ensure interpretability for scientific discovery. These are nontrivial problems beyond the reach of existing skill- or motor-primitive approaches. To our knowledge, this is the first work to introduce offline RL-based imitation learning for behavior segmentation, yielding interpretable skill representations and their compositions.
Appendix J Identifiability Discussion
In this section we clarify the notion of identifiability and how it applies to our motif-based continuous dynamics (MCD) model. A parameterization is said to be identifiable if different parameter values lead to different model predictions. Conversely, if multiple distinct parameterizations yield identical likelihoods or value functions, the model is only identifiable up to an equivalence class (e.g., scaling, rotation, or permutation).
Our model is identifiable at the subspace level.
In MCD, both the transition operator and the action-value function admit the spectral form
$P(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu(s^{\prime}),\qquad Q(s,a)=\phi(s,a)^{\top}u.$
Although individual components of $\phi$, $\mu$, and $u$ may be rescaled by an invertible linear transformation without altering $P$ or $Q$, the subspace spanned by the motif functions is uniquely determined (Ren et al., 2022). This is analogous to PCA, where the principal subspace is identifiable even if the basis vectors within that subspace are not uniquely determined. In practice, they converge well.
Moreover, in the continuous version, the identifiability of the new representations $\psi,\nu$ up to rescaling has been guaranteed by the normalization operation before the output (Appendix C), which also ensures the trainability of the contrastive learning algorithm.
Comparison to HMMs and latent variable models.
Classical behavioral segmentation methods (e.g., Keypoint-Moseq as an HMM variant, SemiSeg as an autoencoder variant) are generally not identifiable without strong additional assumptions. Different transition matrices, emission matrices, or latent dynamics parameters can produce indistinguishable observation distributions. These models typically suffer from label-swapping and more severe transform equivalences in the latent space, meaning that even the latent subspace is not uniquely determined. In contrast, our method fixes the ambient representation through the observed state–action pairs and learns a function class whose span is uniquely tied to the transition operator and value function. Thus, while our motif basis is identifiable only up to linear transforms (as is standard for spectral decompositions), the underlying subspace and its behavioral interpretation are stable, reproducible and identifiable.
Summary.
MCD inherits the identifiability properties of spectral RL methods: the motif basis is identifiable up to an invertible transformation, and the motif subspace—the structure that determines $P$, $Q$, and $\pi$—is uniquely recoverable. This stands in contrast to latent-variable models such as HMMs or SLDS, which in general do not admit identifiable latent subspaces without strong, often unrealistic, constraints.
Appendix K Long-term Dependency Discussion
Compared with prior approaches, our framework implicitly captures the long-term temporal structure of behavior through the Bellman equation in the reinforcement learning (RL) formulation. The component responsible for encoding long-term and multi-scale dynamics is the motif weight $u(t)$. From the linear decomposition of the state–action value function (Eq. 3),
$u(t)=w+\gamma\int V(s^{\prime})\,\mu(s^{\prime})\,q(s^{\prime})\,ds^{\prime},$
it follows that $u(t)$ integrates two sources of information simultaneously: (i) long-horizon structure through the value function $V(s^{\prime})$, which summarizes discounted future trajectories with an infinite horizon, and (ii) short-horizon structure through the immediate reward parameter $w$. Therefore, this formulation naturally enables modeling of multi-scale temporal dependencies. Empirically, we observe this structure in Fig.4B: some motifs exhibit stable, slowly varying activation profiles (e.g., motif 1, gray curve), whereas others fluctuate rapidly and capture short-term transitions (e.g., motif 2, pink curve). Intuitively, consider a mouse that intermittently sniffs throughout the day but engages in fast running only during a brief morning period. The coefficient $u_{i}(t)$ for sniffing would rise gradually across long time windows, while the coefficient $u_{j}(t)$ for running would increase only during short, specific intervals.
Appendix L Hyperparameter Discussion
L.1 Motif dimension
To evaluate the sensitivity of our model to the dimension of the hidden state, we change the dimensions of motifs $d$ and verify them in different datasets.
In the discrete case, there exists a minimum dimension that could fully represent the motif space. In fact, the motif discovery (Eq. 2) is equivalent to SVD of the transition matrix $P(s^{\prime}|s,a)$(Ren et al., 2022). Since the rank of the transition matrix is $|\mathcal{S}|$, the ranks of $\phi$ and $\mu$ matrices are both $|\mathcal{S}|$, which is exactly the minimum dimension of $\phi(s,a)$ that could fully represent the motif space. This could be clearly seen in the gridworld dataset (Fig. 8) and labyrinth navigation dataset (Fig. 9).
For the gridworld dataset, $d_{min}=3*3=9$. If $d<9$, the SVD could only provide a low-rank approximation. As a result, the recovered reward map is blurred and incorrect (Fig. 8A-C) and the correlation remains low (Fig. 8G). For $d\geq 9$, the motif matrix is full-rank. So the reward maps are nearly perfect (Fig. 8D-F), and the correlations are near 100% (Fig. 8G). Similar analysis also applies to labyrinth navigation dataset where $d_{min}=127$(Fig. 9). In this case, variant with fewer motif dimensions ($d=64$) cannot generate realistic reward maps while variants with more motif dimensions ($d\geq 127$) can faithfully recover the reward peaks in each task.
Figure 8: Reward maps generated by models of different motif dimension $d$ in the gridworld dataset.
Figure 9: Reward maps generated by models of different motif dimension $d$ in the labyrinth navigation dataset.
In the continuous case, since the transition kernel is infinite-dimension, any finite $d$ would only provide a low-rank approximation to the transition kernel. As $d$ increases, the fitting performance would be better. However, since the importance of different motifs is different, by observing the performance under different $d$, we could select one to be as small as possible while still capturing the entire motif space as comprehensively as possible, and maintain the performance. In this experiment (Fig. 10), we select $d=64$.
Figure 10: AUC generated by models of different motif dimension $d$ in the animal free-moving dataset.
L.2 Noise distribution
The continuous version of MCD is based on contrastive learning. Therefore, a high-quality negative sample distribution is crucial for the motif and policy learning. Here we evaluate different selections of noise distributions in the continuous free-moving animal behavior dataset.
For the motif learning (Eq. 7), in the main text, we directly sample states from the dataset. In other words, the noise distribution is $\rho(s)=\int\tau^{e}(s,a)da$.
In ”Motif-N” variant, we replace it by a uniform distribution, whose dimension-wise bounds are determined as the maximum and minimum values of the states in the dataset, i.e. $s^{\prime}_{i}\sim U(\min(\rho(s_{i})),\max(\rho(s_{i})))$. Because the negative samples are not good enough, the resulting energy-based model can neither capture high-quality motifs nor provide good basis vectors for later policy learning. Therefore the AUC score remains low (Fig 11).
For the policy learning (Eq. 8), in the main text, we also sample negative samples directly from the dataset $(s^{\prime},a^{\prime})\sim\tau^{e}$ and use $a^{\prime}$ as a negative sample of actions. In other words, the noise distribution is $\zeta(a)=\int\tau^{e}(s,a)ds$.
In the ”Policy-N” variant, we sample negative pairs from uniform distribution $a^{\prime}_{i}\sim U(\min(\zeta(a_{i})),\max(\zeta(a_{i})))$, similar to ”Motif-N”. In the ”AR-Actor” variant, we additionally train an auxilliary autoregressive actor $\pi(a|s)$ to fit $\pi^{e}(a|s)$ using MLE. And the negative samples are from this actor $\pi(a|s)$. For each dimension, the actor has a network of two hidden layers, with hidden dimension=256 each, to produce the mean and variance of this action dimension, i.e. $a_{i}\sim\mathcal{N}(\mu(a_{i}|s,a_{<i}),\sigma^{2}(a_{i}|s,a_{<i}))$. As expected, uniform noise distribution is not good enough, and ”Policy-N” achieves lower AUC than the original version. And the ”AR-Actor” achieves similar performance on the test dataset, proving the robustness of MCD to noise distribution if it is good enough.
Figure 11: AUC generated by models trained under different noise distribution in the animal free-moving dataset.
We also tested the influences of the number of negative samples (Fig. 12). MCD is consistently better than Key-Moseq and maintains a stable performance. It turns out that the model is robust against the choice of number of negative samples.
Figure 12: AUC generated by models trained using different numbers of negative samples for motif contrastive learning in the animal free-moving dataset. $nn=k$ means the number of negative samples per positive sample is $k$.
L.3 Dataset Partition
In the third experiment, we randomly split the whole dataset into training set and test set. Here, we train the model on some mice and then test it on other heldout mice, i.e. split the dataset by mouse identities. This experiment shows whether the motifs learned by the model are general and transferrable across different subjects. The result is shown in Fig. 13 where the complete exclusion of one animal from the training set does not impair the performance. This experiment proves that the MCD can learn a general set of motifs that could be transferred across animals.
Figure 13: AUC generated by models trained using different train-test split methods. For MCD(random), the dataset is randomly split. For MCD(heldout), the behavior of one mouse is specially held out as the test set while the rest is taken as the training set.
L.4 Gaussian random walk prior
In the free-moving animal behavior dataset (Sec. 4.3), we impose a Gaussian random-walk prior on the time-varying motif weights $u(t)$ to encourage temporal smoothness. To assess the sensitivity of our method to this regularization, we varied the strength of the Gaussian prior and re-evaluated model performance (Fig. 14). As expected, very large coefficients introduce slight degradation due to oversmoothing. However, across all tested values, performance remains substantially higher than the baseline model. This demonstrates that MCD is robust to the choice of Gaussian random-walk regularization strength and does not rely on fine-tuning this hyperparameter to achieve strong results.
Figure 14: AUC generated by models tested using different Gaussian random walk prior coefficients in the animal free-moving dataset.
Appendix M Scientific Impact and Relations to Neuroscience and Ethology
Scientific relevance to neuroscience.
Our work is primarily motivated by the neuroscience perspective, where the goal is to extract interpretable behavioral structure that can be directly linked to neural circuit dynamics, internal motivational states, and decision-making processes. From this viewpoint, MCD provides scientifically meaningful variables: it discovers low-level motor motifs that correspond to reproducible movement primitives and models behavior as smoothly varying mixtures of these primitives. This mirrors known neural control principles in the motor cortex, basal ganglia, and brainstem, where overlapping action components, not discrete switches, combine to generate natural movement. The resulting motif representations and time-varying policy weights offer a rich, biologically interpretable representation for analyzing how neural populations evolve alongside behavior.
The learned representations can be mapped to neural representations in future work. Each motif $\phi(s,a)$ defines a low-level motor primitive, giving a clear behavioral regressor for examining whether neurons in motor cortex encode specific movement components. The time-varying policy weights $w(t)$ and reward-related weights $u(t)$ describe how motifs are combined and modulated over time, offering hypotheses about potential control- and value-related signals in the dorsal striatum. While we do not perform neural analyses in this paper, the structured motif representation provides a principled framework for relating behavior to activity in these circuits in future neuroscience studies.
For the training protocol, our method is trained purely from behavior; no neural signals, reward labels, or joint behavior–neural objectives are used. This distinguishes our approach from models that rely on neural data to infer latent goals or value functions. On the other hand, although our method does not require neural recordings, the learned motifs correspond to low-dimensional dynamical components of behavior that can in principle be aligned with neural manifolds (e.g., cyclic modes for gait, motor primitives, or population attractors).
Relevance to ethology and real-world behavior.
Beyond neuroscience, MCD is well suited for ethological studies where the goal is to characterize behavior in naturalistic, minimally constrained environments across multiple timescales. By capturing long-range dependencies, allowing motifs to co-occur, and modeling continuous dynamics, MCD can describe multi-scale organization of behavior, such as exploratory sequences, foraging patterns, or grooming hierarchies, without assuming discrete states. This enables ethologists to quantify how natural behaviors are composed, how they transition, and how they evolve over long durations. Thus, while our emphasis is on neuroscience applications, the method is fully compatible with ethological framework and supports general scientific questions about the structure and function of natural behavior.
Appendix N Generated Rollout Trajectories
In the animal free-moving dataset, we train a generative energy model to estimate the state–action value function. Here, we assess its reliability using Hamiltonian Monte Carlo (HMC) sampling. For each motif $d_{i}$, we construct a one-hot vector $u$ with $u_{d_{i}}=1$ and $u_{d_{j}}=0$ for all $j\neq i$, and then sample an action $a$ from the learned energy model. For each $d_{i}$, we sample $10$ actions in total. The resulting motif-specific rollout trajectories (Fig 15) largely align with the empirical averages (Fig. 7B), demonstrating that the model captures the dominant behavioral tendencies. A small number of trajectories deviate from the empirical trends, probably because $u$ is never used as a strict one-hot vector during training but instead appears as a mixture over motifs. We attribute these deviations to unavoidable generative noise and the distributional mismatch between test-time one-hot inputs and the mixed representations observed in real data.
For that, we choose several clips and generate the rollouts (number of rollout steps=5) based on the inferred $u(t)$. The alignment between generated trajectories and real trajectories (Fig. 16) show that MCD could capture the real behavior dynamics.
HMC parameters.
We use a step size of $1\times 10^{-4}$, 100 leapfrog steps per action, temperature =1=1, mass matrix $M=I$, and initial momentum $r\sim\mathcal{N}(0,1/200)$. The leapfrog integrator is employed to improve sampling accuracy.
Figure 15: Motif-specific rollout trajectories generated by HMC sampling from the trained energy model. Indexes are consistent with Fig. 7.
Figure 16: Rollout trajectories based on inferred $u(t)$ generated by HMC sampling from the trained energy model.